### v1

# LocalSearchBench: Benchmarking Agentic Search in Real-World Local Life Services

Abstract.
Recent advances in large reasoning models (LRMs) have enabled agentic search systems to perform complex multi-step reasoning across multiple sources. However, most studies focus on general information retrieval and rarely explores vertical domains with unique challenges. In this work, we focus on local life services and introduce LocalSearchBench, which encompass diverse and complex business scenarios. Real-world queries in this domain are often ambiguous and require multi-hop reasoning across merchants and products, remaining challenging and not fully addressed. As the first comprehensive benchmark for agentic search in local life services, LocalSearchBench includes over 150,000 high-quality entries from various cities and business types. We construct 300 multi-hop QA tasks based on real user queries, challenging agents to understand questions and retrieve information in multiple steps. We also developed LocalPlayground, a unified environment integrating multiple tools for agent interaction. Experiments show that even state-of-the-art LRMs struggle on LocalSearchBench: the best model (DeepSeek-V3.1) achieves only 34.34% correctness, and most models have issues with completeness (average 77.33%) and faithfulness (average 61.99%). This highlights the need for specialized benchmarks and domain-specific agent training in local life services. Code, Benchmark, and Leaderboard are available at localsearchbench.github.io
Agentic Search, Large Reasoning Model, Local life services
*Project leader.
[…]
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2026/06copyright: noneccs: Computing methodologies Natural language processingccs: Computing methodologies Artificial Intellegence
1. Introduction
(a) General multi-hop QA
(b) Local life service single-hop QA
(c) Local life service multi-hop QA
Figure 1. Illustration of different types of QA tasks
Recent advances in large reasoning models (LRMs) have significantly enhanced the capabilities of AI agents, enabling them to perform complex reasoning and planning tasks (xu2025towards; ferrag2025llm; wu2025agentic; li2025webthinker; ke2025survey; chai2025rlfactory). This progress has driven the development of agentic search systems, which can decompose complex queries, execute multi-step reasoning, and dynamically orchestrate multiple information sources to achieve autonomous cross-domain reasoning (sun2025simpledeepsearch; zhang2025evolvesearch; shi2025pangudeepdiveradaptivesearch; shi2025toollearningwildempowering; zhang2025web; li2025search; mo2025conversational; yue2025promoting; yue2025uiorchestra). These systems also demonstrate the potential to expand from general domains to vertical application scenarios such as e-commerce (zhou2024webarenarealistic; yao2023webshopscalable). However, there has been little work applying agentic search to the local life service domain, and the lack of suitable evaluation benchmarks further constrains new research in this area.
Local life services cover a wide range of sub-domains, including dining, lifestyle, shopping, accommodation, travel, and healthcare (lan2025benchmarking; 10.1145/3580305.3599874). Leveraging location-based digital platforms, these services connect users with nearby merchants and service providers, forming complex networks of businesses, user preferences, and location-dependent service offerings, and creating a vast market (10.1145/3534678.3539122; kusk2025flexible; liu2025mrgrp). As shown in Figure 1(1(c)), real-world user queries often involve multi-constraint merchant recommendation, spatiotemporal service chain planning, and event-driven service bundling, which require complex multi-step processes to deliver the final response. lan2025benchmarking target at this domain and propose the first systematic framework, but is limited to single-hop queries and cannot fully reflect the challenges of real business scenarios. Therefore, there is an urgent need for a multi-hop benchmark tailored to local life scenarios, accurately reflecting the complexity of real-world search tasks and fully challenging existing agentic search methods.
In this paper, we propose an agentic search framework LocalPlayground to evaluate LRMs for local life service. Given the absence of appropriate benchmarks and the critical need for comprehensive evaluation systems, our approach addresses these challenges through comprehensive dataset construction, systematic evaluation framework and targeted model evaluation.
We introduce LocalSearchBench, a comprehensive benchmark comprising over 150,000 data entries across 6 categories: dining, shopping, accommodation, travel, and healthcare, spanning 3 major cities in China. These entries are transformed from raw data into high-quality augmented and anonymized datasets.
Based on these data, we propose a high-quality instruction synthesis method to design 300 multi-hop QA tasks across three reasoning dimensions: multi-constraint merchant recommendation, spatiotemporal service chain planning, and event-driven service bundling. These multi-hop QA tasks are constructed through a multi-agent data generation process, requiring agents to perform complex reasoning across multiple information sources and reasoning steps.
We build a comprehensive evaluation environment with RAG based on LocalSearchBench data and web search tools to simulate real-world local life service search scenarios where models need to retrieve and synthesize information from multiple sources, and conduct evaluations of multiple LRMs.
In summary, our contributions are threefold.
  • To the best of our knowledge, we are the first to build a comprehensive offline high-quality real-world agentic search benchmark for local life service.
  • We propose a systematic agentic search framework LocalPlayground and tool environment for evaluating and applying LRMs in local life services.
  • We evaluate existing models and find that untrained models struggle to complete these tasks, while trained models perform poorly on domain-specific professional requirements.
2. Related Work
2.1. Agentic Search
gou2025mind2web2evaluatingagentic define agentic search as systems capable of integrating multiple auxiliary tools to autonomously perform interactive generation through search tools (e.g., web search APIs or external databases) for handling complex search tasks. Leveraging the powerful language understanding capabilities of LLMs, these search agents can proactively address ambiguous or underspecified queries through clarification or reasonable assumptions (yang2025agentic). They decompose the initial search task into manageable sub-queries, perform dynamic reasoning and planning based on accumulated contextual information, adjust search strategies in real time, and synthesize information from multiple sources (li2025towards).
Recent advances in the reasoning abilities of LRMs further foster the development of LRM-based agentic search systems. The most representative products include OpenAI Deep Research (jaech2024openai), Gemini Deep Research (comanici2025gemini), and Tongyi Deep Research (tongyidr). In addition to accessing real-time information through search APIs, these systems leverage various auxiliary tools to enable deeper and broader research on highly complex problems. In the open-source community,  li2025search first proposed integrating the agentic search workflow into the o1-style reasoning process of LRMs to achieve autonomous knowledge supplementation.  li2025webthinker,  jin2025search, and  song2025r1 further enhanced the autonomous search capabilities of LLMs by applying reinforcement learning, enabling the models to generate queries and invoke external retrieval systems during reasoning.
These methods have made significant progress in general information retrieval and question answering. However, there has been little advancement in vertical domains, especially in areas related to local lifestyle services, due to the challenges associated with training domain-specific agents and the lack of high-quality evaluation benchmarks. In this paper, we propose LocalSearchBench, a comprehensive benchmark specifically designed for evaluating LRMs in local life service scenarios.
2.2. Benchmarking Agentic capabilities
Existing benchmarks for agentic search systems in general domain can be broadly categorized into closed-ended and open-ended QA tasks. Closed-ended QA benchmarks include multi-hop reasoning datasets that require synthesizing information from multiple sources (yang2018hotpotqa; ho2020constructing; press2022measuring; trivedi2022musique), challenging QA benchmarks designed for long-horizon questions with long-tail knowledge (wei2025browsecomps; zhou2025browsecompzhbenchmarking; xi2025infodeepseek; pham2025sealqa), and fact-checking tasks that evaluate claim verification abilities (wei2024long; wadden2020fact; jiang2020hover; wang2024mfc). Open-ended QA benchmarks focus on deep information seeking tasks, targeting multi-perspective queries (rosset2024researchy) and expert-level research tasks (du2025deepresearch; bosse2025deep; tan2024proxyqa), with some incorporating multi-modal capabilities (yuan2025videodeepresearch).
Despite covering various reasoning abilities, existing benchmarks differ from real user queries in local life services and lack assessment of multi-domain coordination. Users often face complex tasks requiring simultaneous service bookings and constraint satisfaction, which current datasets do not adequately evaluate.
However, although these benchmarks cover a range of reasoning capabilities, they have significantly different distribution from real user queries in local life service scenarios, lacking assessment of the multi-domain service coordination required in real-world applications. In reality, users often need to accomplish complex tasks that require coordinating multiple services, such as simultaneously booking restaurants, arranging transportation, and planning entertainment, all while satisfying geographic and temporal constraints. Existing specialized datasets are insufficient for evaluating challenging scenarios like event planning that require multiple API calls across different service domains and real-time constraint satisfaction.
We propose LocalSearchBench to assess multi-hop reasoning and coordinated retrieval for complex local life service tasks, and LocalPlayground as evaluation frameworks that comprehensively measure performance in these specialized scenarios.
3. Construction of LocalSearchBench
Figure 2. Workflow of Local Merchant Database Construction
In this section, we introduce LocalSearchBench, a novel benchmark for agentic search in local life services. LocalSearchBench consists of two main components: Local Merchant Database (Section 3.1) and a multi-hop QA benchmark (Section 3.3).
3.1. Local Merchant Database Construction
In this subsection, we present the construction of the local merchant database for LocalSearchBench, which serves as the foundation for multi-hop QA tasks. As shown in Figure 2, our database construction consists of two stages: Merchant Seed Data Collection and Data Augmentation and Synthesis.
3.1.1. Merchant Seed Data Collection
To ensure LocalSearchBench reflects real-world local life service complexities, we built our dataset based on source data from a leading local life service platform M. We sampled over 260,000 real merchant records spanning from January 2025 to December 2025, covering diverse service categories and geographical distributions across three major Chinese cities, with the generation of these seed data requiring 114 hours of computation time.
Multi-Scenario Coverage
Based on platform M’s core business verticals, we identified six primary service categories reflecting real-world user demand: Dining (35%), Lifestyle (25%), Shopping (20%), Accommodation (10%), Healthcare (5%), and Tourism (5%), as shown in Figure 11. We curated a keyword pool of 411 unique terms across these categories to reflect authentic user search behaviors and regional preferences.
Multi-City Coverage
The collection spans 3 major Chinese cities with comprehensive coverage of benchmarks (Table 1). The location database includes 859 landmarks across diverse categories: commercial centers, transportation hubs, cultural landmarks, educational institutions, medical facilities, and residential areas. Figure 3 shows word clouds of landmark distributions across the 3 cities.
Table 1. Geographic coverage across 3 major Chinese cities
City Landmarks Districts
Shanghai 295 16
Beijing 287 16
Guangzhou 277 11
Total 859 43
(a) Shanghai
(b) Beijing
(c) Guangzhou
Figure 3. Word clouds showing the landmark locations across 3 major Chinese cities.
Merchant Seed Data Retrieval Query Generation
The system employed scenario-specific keyword pools, constructing queries through random combinations of 1-2 keywords to simulate natural search patterns. This approach generated diverse query variations from the 411-keyword foundation, ensuring comprehensive coverage of local search intents across regional preferences.
This multi-city keyword strategy captured regional variations across 43 administrative divisions, providing robust coverage of China’s diverse local search ecosystems as the foundation for LocalSearchBench database synthesis.
3.1.2. Data Augmentation and Synthesis
Data Augmentation
The original merchant seed data contains only 12 basic fields with many missing or unstandardized attributes, which is inadequate for agentic search tasks that require dense, high-coverage merchant profiles. We employ a Data Augmentation Agent to expand these 12-field records into 29-field profiles using LLMs across six dimensions (see Figure 2 and Figure 12). The agent enriches each merchant’s basic, locational, operational, service, contact, and promotional information through 17 newly added fields, as illustrated in Figure 2. The enhanced dataset covers: (1) basic information with standardized categories and descriptions; (2) location intelligence including structured geographical data, coordinates, and landmarks; (3) business operations such as operating hours and pricing; (4) service characteristics like facility lists; (5) contact information, including phone numbers and emails; and (6) promotional content, such as product descriptions and discount offers.
Privacy Rewriting
To comply with data protection regulations and enable safe public release, we employ a Privacy Rewriting Agent on the augmented data (see Figure 13). Real-world merchant data contains personally identifiable information and business-sensitive details that must be anonymized. The agent rewrites 6 privacy-related fields to anonymize merchant identity, location, and contact details while keeping the remaining business information usable, as illustrated in Figure 2.
3.1.3. Data Validation
A Validation Agent, implemented as an LLM-as-a-Judge framework, conducts a comprehensive, two-stage quality assessment to ensure both data quality and privacy compliance in the generated database. Detailed instructions have been carefully designed for augmented data validation and privacy rewriting validation, as illustrated in Figure 14.
Augmented Data Validation
It evaluates augmented Data quality against original seed data across six core dimensions: (1) Data Completeness (20% weight) assessing field coverage and information richness; (2) Logical Consistency (30% weight) validating business logic coherence and field correlations including name format compliance, category alignment, and pricing rationality; (3) Geographic Accuracy (10% weight) verifying location information correctness and address format consistency; (4) Content Quality (20% weight) evaluating textual descriptions, language expression, and business relevance; (5) Format Standardization (10% weight) ensuring data format compliance and structural consistency; and (6) Innovation Degree (10% weight) measuring content creativity and merchant name diversity through batch-level uniqueness analysis. Each evaluation metric corresponds to specific field comparisons as detailed in Table 2(4).
Privacy Rewriting Data Validation
It evaluates privacy rewriting data quality against augmented data across four core dimensions: (1) Anonymization Degree (50% weight) ensuring complete merchant identity anonymization including name deidentification, location sanitization, and contact information anonymization; (2) Logical Consistency (20% weight) maintaining field relationships and business logic coherence after privacy rewriting; (3) Content Consistency (20% weight) preserving non-sensitive information alignment with previous stage outputs; and (4) Innovation Degree (10% weight) assessing the creativity and diversity of anonymized merchant names to prevent pattern recognition. Each evaluation metric corresponds to specific field comparisons as detailed in Table 2(4).
Table 2. Evaluation Frameworks for Data Quality Assessment
(a) Augmented Data Quality
Evaluation Dimension Weight Evaluation Fields
Data Completeness 20% All
Logical Consistency 30% name, category, subcategory, price_range, business_hours, facilities
Geographic Accuracy 10% city, district, landmark, address
Content Quality 20% description, rating, review_count, tags, promotions, products, group_deals
Format Standardization 10% phone, mobile, email
Innovation Degree 10% name
(b) Privacy Writing Data Quality
Evaluation Dimension Weight Evaluation Fields
Anonymization Degree 50% name, address, email
Logical Consistency 20% description, phone, mobile
Content Consistency 20% category, subcategory, city, district, landmark, price_range, business_hours, rating, review_count, delivery_available, delivery_fee, delivery_range, min_order_amount
Innovation Degree 10% name
3.1.4. Local Merchant Database Statistics
Overall, we curated 150,031 representative samples from a pool of over 260,000 original seed records to create synthesized data, which constitute our LocalSearchBench dataset.
The proportions of each category, the number of data for each city are presented in Figure 4. The distribution closely mirrors that of the source data, ensuring our benchmark maintains the same categorical balance as real-world local life service platforms. The geographical distribution of merchant data for each city was derived through geospatial inference to generate latitude-longitude heatmaps, as shown in Figure 5. The distribution patterns align well with urban clustering phenomena and regional economic hierarchies, demonstrating the dataset’s realistic geographical representation.
(a) City distribution
(b) Shanghai
(c) Beijing
(d) Guangzhou
Figure 4. City distribution and category distribution of 150,031 merchant data across three cities.
(a) Shanghai
(b) Beijing
(c) Guangzhou
Figure 5. Geographical distribution heatmaps of 150,031 merchant data across three cities.
3.2. LocalRAG Construction
To enable efficient and accurate merchant information retrieval for LRMs high-performance tool calling, we further construct a comprehensive Retrieval-Augmented Generation (RAG) system, which we term LocalRAG, specifically designed as a tool interface for LRMs. Built upon our previously generated merchant dataset, LocalRAG consists of three core components: vector embedding generation, similarity retrieval, and result reranking, as shown in Figure 6.
3.2.1. Vector Embedding Generation
LocalRAG employs a state-of-the-art embedding model to convert merchant information into high-dimensional vector representations. The process begins by transforming structured merchant data into coherent natural language descriptions. Each merchant record undergoes a text formatting pipeline that systematically concatenates key fields including name, category, address, and service descriptions into readable text. The complete merchant profile is then directly encoded and stored as indexed vector chunks in the vector database, ensuring the embeddings capture both explicit attributes and semantic relationships for nuanced similarity matching.
3.2.2. Similarity Retrieval
The generated embeddings are organized in an indexed vector database with cosine similarity-based search capabilities for fast approximate nearest neighbor retrieval, implementing dense retrieval methodology. During query processing, user queries are embedded using the same model and matched against the indexed merchant vector chunks to retrieve the top-N semantically similar merchant candidates.
3.2.3. Result Reranking
To improve retrieval accuracy, LocalRAG first retrieves the top-N merchants from the vector database as a larger candidate set, then applies a specialized reranking model to perform fine-grained relevance scoring between the query and these candidates. The reranking model computes precise relevance scores for each candidate merchant, ultimately selecting the top-K most relevant merchants based on these refined scores to generate the final answer for LRMs.
Figure 6. Workflow of LocalRAG
3.3. Multi-hop QA Construction
Figure 7. Workflow of Multi-hop QA Construction.
In this subsection, we present the multi-hop QA construction of LocalSearchBench. Building upon the data construction pipeline described in the previous subsections, our multi-hop QA construction pipeline consists of four main stages: seed question collection, question instantiation, answer collection, and data validation. Each of these stages is detailed in the following subsubsections.
3.3.1. Seed Question Collection
To ensure LocalSearchBench reflects real-world local life service complexities, we conducted a statistical analysis to understand user search behaviors and service interactions. Based on requirement complexity and reasoning demands, we classify questions into five intelligence levels (Table 3): L1 (Basic) for simple factual queries; L2 (Advanced) for ambiguous or multi-condition searches; L3 (Intelligent) for composite requirements with multi-turn reasoning; L4 (Expert) for personalized planning; L5 (AGI) for cross-platform coordination with adaptive reflection.
Table 3. Intelligence Level Classification for Seed Questions.
Intellig- ence Level Grade Requirement Understanding Planning-Search-Reflection Loop
L1 Basic
Standard expression of single
precise requirement
Supply matching search with $\leq$
2 conditions
L2 Advanced
Ambiguous expression of single
precise requirement;
Standard expression of
single ambiguous requirement
Supply matching search with $\geq$
3 conditions;
Semantic retrieval search supply;
Multi-source information
integration supply
L3 Intelligent
Composite requirements;
Multi-turn conversational
requirements
Reasoning with certain complexity;
Semantic retrieval with external
information dependency;
Procedural planning requirements;
Comparison and contrast scenarios
L4 Expert
Personalized requirements
External hard-to-access dependency;
Open-ended planning requirements
L5 AGI
Cross-platform coordination;
Complex decision-making;
Exception handling and reflection
Cross-platform dependency;
Task execution;
Adaptive reflection
We collect 1,200 single-hop questions from local life service platform M (90%) and enriched sources (10%) across six service categories: dining, lifestyle, shopping, accommodation, healthcare, and tourism. The platform data captures authentic user queries, while enriched data provides coverage for complex scenarios underrepresented in natural logs. Following the intelligence level taxonomy in Table 3, we focus on L3 (composite requirements) and L4 (personalized planning) questions, as they require multi-step reasoning within a static environment, while L5 tasks involve dynamic cross-platform interactions that are difficult to simulate offline.
(a) Intelligence Level Distribution
(b) Example Questions
Figure 8. Intelligence Level Analysis for Collected Questions
3.3.2. Question Instantiation
We instantiate the collected questions with city-specific information through a two-phase process. First, we construct city profiles for each of the 3 cities (Section 3.1.1), collecting structured information from public resources including landmarks, cultural events, local cuisine, price bands, etc. The collected city data is preprocessed using automated scripts to remove duplicates, normalize formats, and correct inconsistencies, then validated by native annotators to ensure factual accuracy and cultural authenticity.
Second, annotators from platform M manually instantiate questions into concrete multi-hop queries. For each city, approximately 100 questions are selected from the collected 1,200 questions and grounded with city-specific information. To transform single-hop questions into multi-hop queries, annotators enrich questions with 2–4 reasoning segments (2 for 3-hop, 3 for 4-hop, 4 for 5-hop) based on business relevance and logical coherence. These segments include: cross-merchant comparison, schedule planning, chained events, and constraint resolution. We maintain target intelligence-level and hop-count distributions as shown in Figures 9(a) and 9(b), prioritizing underrepresented categories to ensure balanced coverage across service categories and geographic regions (Sections 3.1.1 and 3.1.1).
(a) Intelligence level distribution of QA
(b) Hop count distribution of QA
Figure 9. Distribution of 300 Multi-hop QA
This process generates 300 grounded question instances. For each question, annotators document the hop count, the retrieval target of each hop, and the required search tool (Web Search or LocalRAG) for each hop, enabling downstream validation to enforce hop-by-hop requirements.
3.3.3. Answer Collection
For each instantiated question, we generate its reference answer through a systematic two-step process, leveraging our pre-constructed LocalRAG described in Section 3.2 along with LLM and human expert validation to ensure comprehensive and accurate ground truth answers.
Step 1: Merchant Retrieval. For each question, LocalRAG retrieves relevant local merchant from our comprehensive business database. LocalRAG maintains an indexed vector database of merchant information, where each merchant record is encoded as high-dimensional vector representations capturing merchant names, categories, addresses, and service descriptions along with their semantic relationships. Using semantic similarity matching combined with geographical proximity analysis, the system retrieves relevant business listings, location metadata, and service information that directly relate to the question context. For complex multi-step questions, LocalRAG performs multiple retrieval operations to gather comprehensive merchant information across different categories and locations, providing the necessary contextual foundation for subsequent reasoning steps.
Step 2: Initial Answer Generation. With the retrieved merchant context from LocalRAG, we utilize state-of-the-art LLMs (GPT-5 (openai2025gpt5) and Claude-4.1-opus (anthropic2025claude41)), assisted by domain experts using collaborative tools, to generate preliminary answers. This process is guided by prompts that emphasize factual accuracy, completeness, and strict adherence to the provided merchant information. The prompts explicitly direct the models to cite specific merchant details and avoid introducing information not present in the retrieved context, thus minimizing hallucination. The resulting question-answer pairs serve as input for the subsequent validation and refinement stage.
3.3.4. Data Validation
To ensure the highest quality and reliability of our benchmark, all preliminary QA pairs undergo a rigorous data validation process beyond the initial generation and refinement described in Section 3.3.3. This section outlines the final quality assurance process designed to meet benchmark-level standards, as illustrated in Figure 7. This multi-phase procedure is conducted by experienced technical specialists with deep domain expertise in local life services and involves both answer revision and quality filtering.
Step 1: Expert Answer Revision. Each generated answer is manually reviewed by our expert team consisting of three paper authors working independently. With the assistance of AI-powered development tools, each reviewer independently verifies the factual accuracy of every claim, assesses the completeness of the explanation, and refines the language for clarity and precision. This human-in-the-loop approach allows for nuanced corrections that automated systems might miss, ensuring each answer is not only correct but also comprehensible and contextually appropriate.
Step 2: Quality Filtering. After revision, the QA pairs are subjected to a final filtering step. QA pairs are retained only if they meet all of the following quality standards: (1) questions are clear, well-formulated, and unambiguous; (2) answers are factually correct and complete after revision; and (3) answers can be sufficiently grounded in real-world local service scenarios and business logic. We also enforce balanced distribution across the service categories and geographic regions defined in Section 3.1.1. This stringent filtering ensures that only the most clear, correct, and valuable QA pairs are included in the final benchmark.
3.3.5. Statistics of Multi-hop QA
Figure 9 summarizes the QA pairs, whose intelligence-level and hop-count distributions align with Figures 9(a) and 9(b).
4. LocalPlayground
To verify the effectiveness of the proposed benchmark, we design LocalPlayground, an agent evaluation framework that employs autonomous agents to assess LLM performance, as shown in Figure 10. The framework consists of two key components that work collaboratively: a Search Agent that retrieves and synthesizes information, and a Validation Agent that evaluates response quality against ground truth.
Search Agent
The Search Agent retrieves relevant information using two search tools: LocalRAG and Web Search.
LocalRAG performs semantic similarity matching and geographical proximity analysis on vector-indexed merchant databases to retrieve business listings, points of interest and service descriptions.
Web Search queries real-time contextual information such as temporal facts, weather conditions, general knowledge, news and events and supplementary product details that may affect merchant operations and user decisions through web requests to search APIs, which interface with search engines (specifically Baidu Search (liu2021pre) in our experiments) to retrieve up-to-date web content.
The Search Agent operates iteratively over N rounds. In each round, it invokes LocalRAG at most once and Web Search at most once, allowing the agent to progressively refine its information gathering strategy based on intermediate results.
Validation Agent
To ensure comprehensive and objective assessment of generated outputs, we employ a Validation Agent implemented as an LLM-as-a-Judge framework, as shown in Figure 10. Given the generated outputs and the golden answer, an LLM (specifically Claude-Sonnet-4 (anthropic2025claude4) in our experiments) evaluates answer quality across seven dimensions: (1) Correctness: measures whether the answer is factually accurate; (2) Completeness: measures whether the answer fully addresses the question, (3) Fluency: measures the readability and coherence of the response; (4) Faithfulness: measures the degree to which the response avoids fabricated or hallucinated content; (5) Safety: measures whether the response avoids harmful or inappropriate content; (6) Avg. tool calls: measures the average number of tool invocations (LocalRAG and web search) and (7) Avg. rounds: measures the average number of conversation rounds.
Dimension (1) employs a binary scoring scheme (0 or 1), clearly distinguishing correct from incorrect responses. For dimensions (2)-(5), each dimension is evaluated using a granular 10-point rating system (0-10), with higher scores indicating superior performance. Scores across dimensions (1)-(5) are normalized and expressed as percentages for standardized comparison. For dimensions Dimensions (6)-(7) capture continuous variables representing frequency counts. To ensure reliability, each instance is evaluated five times per dimension, with final scores computed by averaging across trials. Candidate systems are anonymized, question order is randomized, and the judge model is separate from all evaluated models to prevent self-assessment.
Figure 10. Overview of LocalPlayground
5. Evaluation
We aim to answer the following research questions:
RQ1 (Benchmark Quality): How well does LocalSearchBench capture the diversity and complexity of real-world merchant search scenarios?
RQ2 (Benchmark Results): How do existing open-source and closed-source LLMs perform on our benchmark?
RQ3 (Sensitivity Study): How does the maximum number of conversation rounds affect search agent performance?
5.1. Evaluated Models
We evaluate ten widely recognized LRMs using LocalPlayground, including five closed-source LRMs, i.e., GPT-4.1 (openai2025gpt4-1), Gemini-2.5-Pro (comanici2025gemini), Qwen-Plus, Hunyuan-T1 (team2025hunyuan), and LongCat-Large-32K (team2025longcat) as well as five open-source LRMs, i.e., DeepSeek-V3.1 (deepseek2025v3), GLM-4.5 (zeng2025glm), Qwen3-235B-A22B, Qwen3-32B, and Qwen3-14B (yang2025qwen3).
Unless specified, LocalPlayground is configured with default settings: (1) a maximum of 5 (N = 5) reasoning-action iterations per question; (2) for LocalRAG, Qwen3-Embedding-8B (zhang2025qwen3) as the embedding model and Qwen3-Reranker-8B (zhang2025qwen3) as the reranking model, with dense retrieval returning top-N = 100 candidates and reranking returning top-K = 20 merchants; (3) for web search, top-k retrieval where k = 20; and (4) LRMs decoding temperature set to 0 to ensure reproducibility. All experiments were conducted on a system equipped with an Intel Xeon Gold 5218 CPU @ 2.30GHz and 1 $\times$ H20-141G GPU.
5.2. Benchmark Quality (RQ1)
To evaluate the quality of our LocalSearchBench database (Section 3.1.1), we employ the two-stage validation framework described in Section 3.1.3. We randomly sample 1,000 merchant entries and adopt an LLM-as-judge approach using Claude-Sonnet-4 (anthropic2025claude4) to assess data quality across multiple dimensions. As shown in Table 4, both augmented and privacy rewriting data demonstrate high quality with overall scores of 0.8596 and 0.9217, respectively.
To further validate the practical utility of our benchmark and mitigate potential biases inherent in automated evaluation, we invited four independent evaluators with expertise in local life service domains, who each evaluated outputs from all methods to ensure comprehensive coverage. They assessed outputs in a blind manner based on data consistency, completeness, and anonymization effectiveness using a 5-point Likert scale (5-Strongly Agree to 1-Strongly Disagree). For analysis, ratings of 4-5 were converted to 1 (satisfied) and 1-3 to 0 (not satisfied).
As shown in Table 7, the human evaluation achieves 86.51% agreement with automated LLM assessment across all 20 fields, validating the effectiveness of our LLM-as-judge approach.
To ensure the quality of our Multi-hop QA, we applied stringent manual verification (Section 3.3.4), retaining only questions with clear, answerable information from the given merchant data and filtering out ambiguous or unanswerable queries.
Finding 1: Automated LLM evaluation shows high quality across LocalSearchBench, and human verification on sampled data confirms the reliability of automated assessment, validating the benchmark’s practical utility.
Table 4. Evaluation Frameworks for Data Quality Assessment
(a) Augmented Data Quality
Dimension Weight Score
Data Completeness 20% 0.8605
Logical Consistency 30% 0.9000
Geographic Accuracy 10% 0.8568
Content Quality 20% 0.8521
Format Standardization 10% 0.9060
Innovation Level 10% 0.7080
Overall 100% 0.8596
(b) Privacy Rewriting Data Quality
Dimension Weight Score
Anonymization Level 50% 0.9473
Logical Consistency 20% 0.8383
Content Consistency 20% 0.9919
Innovation Level 10% 0.8200
Overall 100% 0.9217
5.3. Benchmark Results (RQ2)
Impact of web search.
We first analyze the impact of web search, comparing LocalPlayground (LRM+ LocalRAG +web search) against LRM + LocalRAG without websearch as the baseline to assess web search effectiveness. We exclude configurations of the base LRM alone or LRM without LocalRAG from our comparison, as multi-hop questions in our benchmark require information exclusively available in the LocalRAG knowledge base.
As shown in Table 5, incorporating web search into LocalPlayground presents a trade-off between answer quality and faithfulness. On average, LocalPlayground improves Correctness by 4.37pp (from 25.58% to 29.95%) and Completeness by 3.95pp (from 73.38% to 77.33%), but reduces Faithfulness by 3.64pp (from 65.63% to 61.99%). This trade-off is most pronounced for DeepSeek-V3.1, which gains 4.57pp in Correctness but loses 8.45pp in Faithfulness. These results suggest that while web search is essential for retrieving critical information in multi-hop reasoning, it also introduces noise that compromises response faithfulness.
Performance across language models.
The choice of underlying LRM significantly influences performance across all metrics. Among all models, DeepSeek-V3.1 achieves the highest Correctness (34.34%), followed by GLM-4.5 (33.78%) and LongChat-Large-32K (33.19%). Qwen-Plus-Latest leads in Completeness (80.94%), while Gemini-2.5-Pro excels in Faithfulness (78.26%). Closed-source models outperform open-source ones in completeness (79.08% vs. 75.57%) and efficiency (2.25 vs. 2.61 tool calls), though differences in Correctness are more modest. Despite these variations, overall performance remains limited, as even the best model achieves only 34.34% Correctness, underscoring the substantial challenge of multi-hop local search.
Table 5. Performance comparison of different models on LocalSearchBench (Max Round N = 5).
Model Config Evaluation Metrics
Avg. tool calls Avg. rounds Correctness (%) Completeness (%) Fluency (%) Faithfulness (%) Safety (%)
Closed-source LRMs
[Uncaptioned image] GPT-4.1 w/o web search 1.05 2.06 18.73 70.64 72.37 75.10 83.71
LocalPlayground 1.72 ($\uparrow$0.67) 2.70 ($\uparrow$0.64) 26.76 ($\uparrow$8.03) 75.42 ($\uparrow$4.78) 71.54 ($\downarrow$0.83) 72.63 ($\downarrow$2.47) 81.00 ($\downarrow$2.71)
[Uncaptioned image] Gemini-2.5-Pro w/o web search 1.32 2.29 22.56 70.51 74.48 75.52 84.51
LocalPlayground 1.89 ($\uparrow$0.57) 2.86 ($\uparrow$0.57) 26.09 ($\uparrow$3.53) 77.93 ($\uparrow$7.42) 73.50 ($\downarrow$0.98) 78.26 ($\uparrow$2.74) 83.81 ($\downarrow$0.70)
[Uncaptioned image] Qwen-Plus-Latest w/o web search 1.79 2.79 27.76 77.19 73.21 68.78 83.41
LocalPlayground 2.59 ($\uparrow$0.80) 3.12 ($\uparrow$0.33) 32.79 ($\uparrow$5.03) 80.94 ($\uparrow$3.75) 73.65 ($\uparrow$0.44) 68.68 ($\downarrow$0.10) 83.14 ($\downarrow$0.27)
[Uncaptioned image] LongCat-Large-32K w/o web search 1.59 2.49 30.93 79.45 71.82 65.58 80.72
LocalPlayground 2.73 ($\uparrow$1.14) 3.22 ($\uparrow$0.73) 33.19 ($\uparrow$2.26) 80.51 ($\uparrow$1.06) 73.40 ($\uparrow$1.58) 60.80 ($\downarrow$4.78) 81.14 ($\uparrow$0.42)
[Uncaptioned image] Hunyuan-T1 w/o web search 1.82 2.65 30.10 74.50 73.53 71.76 84.50
LocalPlayground 2.30 ($\uparrow$0.48) 3.15 ($\uparrow$0.50) 32.94 ($\uparrow$2.84) 80.59 ($\uparrow$6.09) 73.00 ($\downarrow$0.53) 63.77 ($\downarrow$7.99) 84.71 ($\uparrow$0.21)
Open-source LRMs
[Uncaptioned image] Qwen3-14B w/o web search 1.20 2.16 21.07 71.54 70.67 54.47 82.71
LocalPlayground 2.04 ($\uparrow$0.84) 2.85 ($\uparrow$0.69) 26.42 ($\uparrow$5.35) 76.25 ($\uparrow$4.71) 70.97 ($\uparrow$0.30) 48.54 ($\downarrow$5.93) 81.54 ($\downarrow$1.17)
[Uncaptioned image] Qwen3-32B w/o web search 1.33 2.10 22.74 68.23 70.40 47.73 80.80
LocalPlayground 2.48 ($\uparrow$1.15) 2.89 ($\uparrow$0.79) 24.75 ($\uparrow$2.01) 71.37 ($\uparrow$3.14) 70.17 ($\downarrow$0.23) 40.84 ($\downarrow$6.89) 81.34 ($\uparrow$0.54)
[Uncaptioned image] Qwen3-235B-A22B w/o web search 1.26 2.10 23.08 66.56 70.50 54.94 81.20
LocalPlayground 2.35 ($\uparrow$1.09) 2.81 ($\uparrow$0.71) 28.43 ($\uparrow$5.35) 73.48 ($\uparrow$6.92) 70.54 ($\uparrow$0.04) 52.43 ($\downarrow$2.51) 81.67 ($\uparrow$0.47)
[Uncaptioned image] GLM-4.5 w/o web search 1.92 2.92 29.10 76.56 67.86 73.15 80.97
LocalPlayground 2.73 ($\uparrow$0.81) 3.66 ($\uparrow$0.74) 33.78 ($\uparrow$4.68) 76.76 ($\uparrow$0.20) 71.81 ($\uparrow$3.95) 73.12 ($\downarrow$0.03) 76.66 ($\downarrow$4.31)
[Uncaptioned image] Deepseek-V3.1 w/o web search 2.36 3.36 29.77 78.64 72.17 69.25 81.56
LocalPlayground 3.43 ($\uparrow$1.07) 4.02 ($\uparrow$0.66) 34.34 ($\uparrow$4.57) 80.00 ($\uparrow$1.36) 72.39 ($\uparrow$0.22) 60.80 ($\downarrow$8.45) 81.48 ($\downarrow$0.08)
Average w/o web search 1.56 2.49 25.58 73.38 71.70 65.63 82.41
LocalPlayground 2.43 ($\uparrow$0.87) 3.13 ($\uparrow$0.64) 29.95 ($\uparrow$4.37) 77.33 ($\uparrow$3.95) 72.10 ($\uparrow$0.40) 61.99 ($\downarrow$3.64) 81.65 ($\downarrow$0.76)
Finding 2: Web search improves correctness and completeness but sacrifices faithfulness. Closed-source models outperform open-source ones in completeness and efficiency. However, overall performance remains low, highlighting the substantial challenge of multi-hop local life service search.
5.4. Sensitivity Study (RQ3)
To investigate the impact of maximum conversation rounds on tool usage and model performance, we conduct a sensitivity study using DeepSeek-V3.1, the best-performing model in correctness from RQ2, with varying max conversation round $N$ from 2 to 8.
As shown in Table 6, the optimal configuration is $N=5$, which achieves the highest Correctness (34.34%). When $N<5$, insufficient rounds limit information gathering, resulting in lower correctness (30.87%–32.11%). When $N>5$, despite increased tool usage, correctness begins to decline (31.32%–32.77%), suggesting that excessive rounds introduce noisy information that hinders answer quality rather than improving it. In terms of computational cost, increasing the maximum rounds initially leads to more tool calls and longer conversations, but the growth saturates beyond $N=5$. The average tool calls plateau around 3.68–3.81 and average rounds stabilize near 4.29–4.48. Therefore, $N=5$ represents a critical turning point that balances performance and efficiency (3.43 tool calls, 4.02 rounds).
Regarding other metrics, Fluency remains stable across all settings (71.44%–72.39%), while faithfulness gradually decreases from 64.78% ($N=2$) to 60.80% ($N=5$), suggesting that longer conversations introduce more opportunities for hallucinations. Safety scores stay consistently high (80.39%–82.39%), indicating robustness across different round limits.
Table 6. Sensitivity analysis of Max Round N on model performance.
Max Round Evaluation Metrics
Avg. tool calls Avg. rounds Correct- ness(%) Complete- ness(%) Fluency (%) Faithful- ness(%) Safety (%)
$N=2$ 1.96 1.98 30.87 79.63 71.44 64.78 81.14
$N=3$ 2.58 2.92 31.12 79.43 71.57 63.50 82.27
$N=4$ 3.20 3.75 32.11 79.70 72.27 64.26 82.39
$N=5$ 3.43 4.02 34.34 80.00 72.39 60.80 81.48
$N=6$ 3.79 4.45 32.77 80.20 71.96 60.56 80.91
$N=7$ 3.81 4.48 31.65 80.20 71.89 61.32 80.84
$N=8$ 3.68 4.29 31.32 79.26 71.65 61.33 80.39
Finding 3: Setting maximum rounds to 5 achieves optimal performance, maximizing correctness while maintaining efficiency. Beyond this threshold, tool usage saturates but correctness declines, indicating diminishing returns from information redundancy.
6. Conclusion
We propose LocalSearchBench, a new benchmark for agentic search in local life services, designed to evaluate multi-hop reasoning in real-world business scenarios. The dataset includes 150,031 samples and 300 multi-hop QA tasks, curated from over 260,000 merchant records in 3 Chinese cities and 6 service categories through data augmentation, privacy rewriting, and rigorous validation. Using the LocalPlayground toolkit for search and seven-dimensional validation, our experiments on ten leading LLMs show that current models still struggle. For example, DeepSeek-V3.1 achieves only 34.34% correctness. By open-sourcing LocalSearchBench and LocalPlayground, we aim to foster progress in domain-specific agentic search and encourage future improvements in data coverage and model training.
Appendix A Appendix
A.1. LocalSearchBench Task Details
Figure 11. Standardized category distribution of real-world data
Table 7. Human-AI Agreement Rate by Field
Field Human-AI Agreement Rate
subcategory 100.00%
district 100.00%
category 100.00%
city 100.00%
review_count 96.97%
price_range 95.96%
business_hours 95.96%
landmark 92.93%
rating 91.92%
delivery_available 91.82%
delivery_range 91.82%
delivery_fee 87.78%
min_order_amount 86.77%
address 77.68%
phone 73.64%
name 72.63%
mobile 71.61%
products 71.61%
email 68.55%
description 62.51%
Average 86.51%
Algorithm 1 Algorithm of Iterative Search
Input:
User query $Q$, Maximum rounds NN
Output: Retrieved information set $\mathcal{I}$
1
// Phase 1: Initialization
$\mathit{information}\leftarrow\emptyset$ ;
// Initialize empty information set
$\mathit{round}\leftarrow 1$ ;
// Initialize iteration counter
2
// Phase 2: Iterative Retrieval Loop
3 for $i\leftarrow 1$ to NN do
4 $\mathit{web\_results}\leftarrow\emptyset$, $\mathit{rag\_results}\leftarrow\emptyset$;
5
 // LLM Response Parsing
6 $\mathit{response}\leftarrow$ LLM($Q$, $\mathit{information}$);
7 $\mathit{web\_query}\leftarrow$ ExtractTag($\mathit{response}$, “<web_search>”);
8 $\mathit{rag\_query}\leftarrow$ ExtractTag($\mathit{response}$, “<rag>”);
9
 // Web Search (at most once per round)
10 if $\mathit{web\_query}\neq\emptyset$ then
11    $\mathit{web\_results}\leftarrow$ WebSearch($\mathit{web\_query}$);
    // Query temporal facts, weather, news, events
12
13   end if
14
 // LocalRAG Retrieval (at most once per round)
15 if $\mathit{rag\_query}\neq\emptyset$ then
16    $\mathit{rag\_results}\leftarrow$ LocalRAGSearch($\mathit{rag\_query}$);
    // Semantic similarity + geographical proximity analysis
    // Retrieve merchants, POIs, service descriptions
17
18   end if
19
20 $\mathit{information}\leftarrow\mathit{information}\cup\mathit{web\_results}\cup\mathit{rag\_results}$;
21
22 if SufficientInfo($\mathit{information}$, $Q$) or $i=N$ then
23    break;
24
25   end if
26
27 end for
28
// Phase 3: Return Results
29 return $\mathit{information}$;
30
Figure 12. Prompt for Data Augmentation
Figure 13. Prompt for Privacy Rewriting
Figure 14. Prompt for Augmentation Validation