### v2

# LocalSearchBench: Benchmarking Agentic Search in Real-World Local Life Services

Abstract.
Recent advances in large reasoning models (LRMs) have enabled agentic search systems to perform complex multi-step reasoning across multiple sources. However, most studies focus on general information retrieval and rarely explores vertical domains with unique challenges. In this work, we focus on local life services and introduce LocalSearchBench, which encompass diverse and complex business scenarios. Real-world queries in this domain are often ambiguous and require multi-hop reasoning across merchants and products, remaining challenging and not fully addressed. As the first comprehensive benchmark for agentic search in local life services, LocalSearchBench comprises a database of over 1.3M merchant entries across 6 service categories and 9 major cities, and 900 multi-hop QA tasks from real user queries that require multi-step reasoning. We also developed LocalPlayground, a unified environment integrating multiple tools for LRMs interaction. Experiments show that even state-of-the-art LRMs struggle on LocalSearchBench: the best model (DeepSeek-V3.2) achieves only 35.60% correctness, and most models have issues with completeness (average 60.32%) and faithfulness (average 30.72%). This highlights the need for specialized benchmarks and domain-specific agent training in local life services. Code, Benchmark, and Leaderboard are available at localsearchbench.github.io.
Agentic Search, Large Reasoning Model, Local life services
*Project leader.
[…]
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2026/06copyright: noneccs: Computing methodologies Natural language processingccs: Computing methodologies Artificial Intellegence
1. Introduction
(a) General multi-hop QA
(b) Local life service single-hop QA
(c) Local life service multi-hop QA
Figure 1. Illustration of different types of QA tasks
Recent advances in large reasoning models (LRMs) have significantly enhanced the capabilities of AI agents, enabling them to perform complex reasoning and planning tasks (Xu et al., 2025; Ferrag et al., 2025; Wu et al., 2025; Li et al., 2025b; Ke et al., 2025; Chai et al., 2025; Chen et al., 2025). This progress has driven the development of agentic search systems, which can decompose complex queries, execute multi-step reasoning, and dynamically orchestrate multiple information sources to achieve autonomous cross-domain reasoning (Sun et al., 2025; Zhang et al., 2025a; Shi et al., 2025a, b; Zhang et al., 2025b; Li et al., 2025a; Mo et al., 2025; Yue et al., 2025b, a). These systems also demonstrate the potential to expand from general domains to vertical application scenarios such as e-commerce (Zhou et al., 2024; Yao et al., 2023). However, there has been little work applying agentic search to the local life service domain, and the lack of suitable evaluation benchmarks further constrains new research in this area.
Local life services are online-offline integrated services connecting local merchants with nearby users, including dining, lifestyle, shopping, accommodation, travel, healthcare and more to meet residents’ daily needs (Lan et al., 2025b, 2023). Leveraging location-based digital platforms, these services form complex networks of businesses, user preferences, and location-dependent service offerings, creating a vast market (Li et al., 2022; Kusk, 2025; Liu et al., 2025a). As shown in Figure 1(1(c)), real-world user queries often involve multi-constraint merchant recommendation, spatiotemporal service chain planning, and event-driven service bundling, which require complex multi-step processes to deliver the final response.
Lan et al. (2025b) target at this domain and propose the first systematic framework, but is limited to single-hop queries and cannot fully reflect the challenges of real business scenarios. Moreover, existing agentic search benchmarks (Table 1) focus on general domains and miss critical local-life challenges such as geographic constraints and multi-service integration. Therefore, there is an urgent need for a multi-hop benchmark tailored to local life scenarios, accurately reflecting the complexity of real-world search tasks and fully challenging existing agentic search methods.
To bridge these gaps, we first introduce LocalSearchBench, a comprehensive benchmark comprising over 1,300,000 data entries across 6 primary service categories, spanning 9 major cities in China. These entries are transformed from raw data into high-quality augmented and anonymized database. Based on these data, we propose a high-quality instruction synthesis method to design 900 multi-hop QA tasks across two intelligence levels. These multi-hop QA tasks are constructed through a multi-agent data generation process, requiring agents to perform complex reasoning across multiple information sources and reasoning steps.
Building upon LocalSearchBench, we further develop LocalPlayground, an agentic search framework to evaluate LLM agents for local life service. Given the absence of appropriate benchmarks and the critical need for comprehensive evaluation systems, our approach addresses these challenges through comprehensive dataset construction, systematic evaluation framework and targeted model evaluation. We build a comprehensive evaluation environment with LocalRAG based on merchant database and web search tools to simulate real-world local life service search scenarios where models need to retrieve and synthesize information from multiple sources, and conduct evaluations of multiple LRMs.
In summary, our contributions are threefold.
  • To the best of our knowledge, we are the first to build a comprehensive offline high-quality real-world agentic search benchmark for local life service.
  • We propose a systematic agentic search framework LocalPlayground and tool environment for evaluating and applying LRMs in local life services.
  • We evaluate existing models and find that untrained models struggle to complete these tasks, while trained models perform poorly on domain-specific professional requirements.
2. Related Work
Table 1. Comparison of existing user interaction benchmarks across key evaluation dimensions. “” indicates fully addressed, “ ” indicates partially addressed, and “” indicates not addressed.
Benchmark Reasoning Complexity Domain Specificity Task Complexity
Multi-hop Reasoning Tool Integration Context Dependency Domain Knowledge Geographic Constraints Multi-service Integration Real-world Scenarios Scale (samples)
HotpotQA (Yang et al., 2018) 113K
MuSiQue (Trivedi et al., 2022) 25K
BrowseComp (Wei et al., 2025) 1K
DeepResearch (Du et al., 2025) 500
DeepWideSearch (Lan et al., 2025a) 220
ResearchY (Rosset et al., 2024) 1.5K
ProxyQA (Tan et al., 2024) 1550
LocalSearchBench (ours) 1.35M + 900
2.1. Agentic Search
Gou et al. (2025) define agentic search as systems capable of integrating multiple auxiliary tools to autonomously perform interactive generation through search tools (e.g., web search APIs or external databases) for handling complex search tasks. Leveraging the powerful language understanding capabilities of LLMs, these search agents can proactively address ambiguous or underspecified queries through clarification or reasonable assumptions (Yang et al., 2025b). They decompose the initial search task into manageable sub-queries, perform dynamic reasoning and planning based on accumulated contextual information, adjust search strategies in real time, and synthesize information from multiple sources (Li et al., 2025c).
Recent advances in the reasoning abilities of LRMs further foster the development of LRM-based agentic search systems. The most representative products include OpenAI Deep Research (Jaech et al., 2024), Gemini Deep Research (Comanici et al., 2025), and Tongyi Deep Research (Team, 2025). In addition to accessing real-time information through search APIs, these systems leverage various auxiliary tools to enable deeper and broader research on highly complex problems. In the open-source community,  Li et al. (2025a) first proposed integrating the agentic search workflow into the o1-style reasoning process of LRMs to achieve autonomous knowledge supplementation.  Li et al. (2025b),  Jin et al. (2025), and  Song et al. (2025) further enhanced the autonomous search capabilities of LLMs by applying reinforcement learning, enabling the models to generate queries and invoke external retrieval systems during reasoning.
These methods have made significant progress in general information retrieval and question answering. However, there has been little advancement in vertical domains, especially in areas related to local lifestyle services, due to the challenges associated with training domain-specific agents and the lack of high-quality evaluation benchmarks. In this paper, we propose LocalSearchBench, a comprehensive benchmark specifically designed for evaluating LRMs in local life service scenarios.
2.2. Benchmarking Agentic capabilities
Existing benchmarks for agentic search systems in general domain can be broadly categorized into closed-ended and open-ended QA tasks. Closed-ended QA benchmarks include multi-hop reasoning datasets that require synthesizing information from multiple sources (Yang et al., 2018; Ho et al., 2020; Press et al., 2022; Trivedi et al., 2022), challenging QA benchmarks designed for long-horizon questions with long-tail knowledge (Wei et al., 2025; Zhou et al., 2025; Xi et al., 2025; Pham et al., 2025), and fact-checking tasks that evaluate claim verification abilities (Wei et al., 2024; Wadden et al., 2020; Jiang et al., 2020; Wang et al., 2024). Open-ended QA benchmarks focus on deep information seeking tasks, targeting multi-perspective queries (Rosset et al., 2024) and expert-level research tasks (Du et al., 2025; Bosse et al., 2025; Tan et al., 2024), with some incorporating multi-modal capabilities (Yuan et al., 2025; Dong et al., 2025; Liu et al., 2025c).
Despite covering various reasoning abilities, existing benchmarks differ from real user queries in local life services and lack assessment of multi-domain coordination. Users often face complex tasks requiring simultaneous service bookings and constraint satisfaction, which current datasets do not adequately evaluate.
However, although these benchmarks cover a range of reasoning capabilities, they have significantly different distribution from real user queries in local life service scenarios, lacking assessment of the multi-domain service coordination required in real-world applications. In reality, users often need to accomplish complex tasks that require coordinating multiple services, such as simultaneously booking restaurants, arranging transportation, and planning entertainment, all while satisfying geographic and temporal constraints. Existing specialized datasets are insufficient for evaluating challenging scenarios like event planning that require multiple API calls across different service domains and real-time constraint satisfaction.
We propose LocalSearchBench to assess multi-hop reasoning and coordinated retrieval for complex local life service tasks, and LocalPlayground as evaluation frameworks that comprehensively measure performance in these specialized scenarios.
3. Construction of LocalSearchBench
Figure 2. Workflow of Local Merchant Database Construction.
In this section, we introduce LocalSearchBench, a novel benchmark for agentic search in local life services. LocalSearchBench consists of two main components: Local Merchant Database (Section 3.1) and a multi-hop QA benchmark (Section 3.3).
3.1. Local Merchant Database Construction
In this subsection, we present the construction of the local merchant database for LocalSearchBench, which serves as the foundation for multi-hop QA tasks. As shown in Figure 2, our database construction consists of three stages: Merchant Seed Data Collection, Data Augmentation and Synthesis, and Data Validation.
3.1.1. Merchant Seed Data Collection
To ensure LocalSearchBench reflects real-world local life service complexities, we built our local merchant database based on raw data from a leading local life service platform M. We sampled over 1,600,000 real merchant records spanning from January to December 2025, covering diverse service categories and landmarks across 9 major Chinese cities, involving over 9,900,000 retrievals from raw data, with the generation of these seed data requiring 1,418 hours of computation time.
Multi-Scenario Coverage
Based on platform M’s core business verticals, we identified six primary service categories reflecting real-world user demand: Dining (35%), Lifestyle (25%), Shopping (20%), Accommodation (10%), Healthcare (5%), and Tourism (5%), as shown in Figure 4(4(a)). We curated a keyword pool of 927 unique terms across these categories to reflect authentic user search behaviors and regional preferences.
Multi-City Coverage
The collection spans 9 major Chinese cities with comprehensive coverage of benchmarks (Table 2). The location database includes 10,741 landmarks across diverse categories: commercial centers, transportation hubs, cultural landmarks, educational institutions, medical facilities, and residential areas. Figure 3 shows word clouds of landmark distributions for three of the nine cities.
Table 2. Geographic coverage across 9 major Chinese cities
City Landmarks Districts Retrievals
Shanghai 1556 16 1,442,412
Beijing 1451 16 1,345,077
Guangzhou 1026 11 951,102
Shenzhen 1039 9 963,153
Hangzhou 1129 13 1,046,583
Chengdu 1073 20 994,671
Chongqing 1008 38 934,416
Wuhan 1040 13 964,080
Suzhou 1149 9 1,065,123
Total 10,741 145 9,956,907
Merchant Seed Data Retrieval Query Generation
We construct scenario-specific queries by randomly combining 1–2 keywords from the 927-term pool to simulate natural search patterns. Applied across 145 administrative divisions, this multi-city keyword strategy generates diverse query variations and comprehensive coverage of local search intents and regional preferences, forming the foundation for LocalSearchBench’s merchant database synthesis.
(a) Shanghai
(b) Beijing
(c) Guangzhou
Figure 3. Word clouds showing the landmark locations across 3 major Chinese cities.
3.1.2. Data Augmentation and Synthesis
Data Augmentation
The original merchant seed data contains only 12 basic fields with many missing or unstandardized attributes, which is inadequate for agentic search tasks that require dense, high-coverage merchant profiles. We employ a Data Augmentation Agent to expand these 12-field records into 29-field profiles using LLMs across six dimensions (see Figure 2 and Figure 14). The agent enriches each merchant’s basic, locational, operational, service, contact, and promotional information through 17 newly added fields, as illustrated in Figure 2.
Privacy Rewriting
To comply with data protection regulations and enable safe public release, we employ a Privacy Rewriting Agent on the augmented data (see Figure 15). Real-world merchant data contains personally identifiable information and business-sensitive details that must be anonymized (Liu et al., 2023; Liu, 2024). The agent rewrites seven privacy-related fields to anonymize merchant identity, location, and contact details while keeping the remaining business information usable, as illustrated in Figure 2.
3.1.3. Data Validation
A Validation Agent, implemented as an LLM-as-a-Judge framework, conducts a comprehensive, two-stage quality assessment to ensure both data quality and privacy compliance in the generated database. Detailed instructions have been carefully designed for augmented data validation and privacy rewriting validation, as illustrated in Figure 14.
Augmented Data Validation
It evaluates augmented data quality against the original seed data along six weighted dimensions: Data Completeness (20%), Logical Consistency (30%), Geographic Accuracy (10%), Content Quality (20%), Format Standardization (10%), and Innovation Degree (10%), as summarized in Table 2(3).
Validation Thresholds: Thresholds are calibrated from an empirical study on 1,000 samples (overall score 0.86), benchmark dataset standards (minimum 0.80), and alignment with human ratings. A merchant record passes if the Overall Score $\geq 0.80$, all six dimensions $\geq 0.70$, key dimensions (Logical Consistency, Data Completeness, Content Quality) meet their target thresholds, and there are no critical failures in essential fields (name format, category alignment, address consistency).
Table 3. Evaluation Frameworks for Data Quality Assessment
(a) Augmented Data Quality
Evaluation Dimension Weight Evaluation Fields
Data Completeness 20% All
Logical Consistency 30% name, category, subcategory, price_range, business_hours, facilities
Geographic Accuracy 10% city, district, landmark, address
Content Quality 20% description, rating, review_count, tags, promotions, products, group_deals
Format Standardization 10% phone, mobile, email
Innovation Degree 10% name
(b) Privacy Writing Data Quality
Evaluation Dimension Weight Evaluation Fields
Anonymization Degree 50% name, address, email
Logical Consistency 20% description, phone, mobile
Content Consistency 20% category, subcategory, city, district, landmark, price_range, business_hours, rating, review_count, delivery_available, delivery_fee, delivery_range, min_order_amount
Innovation Degree 10% name
Privacy Rewriting Data Validation
It evaluates the quality of privacy rewriting against the augmented data along four dimensions: Anonymization Degree (50%), Logical Consistency (20%), Content Consistency (20%), and Innovation Degree (10%), as detailed in Table 2(3).
Validation Thresholds: Due to the higher privacy risk, this stage adopts stricter criteria informed by GDPR requirements and an empirical study on 1,000 samples (overall score 0.92, Anonymization Degree 0.95) (Liu et al., 2025b). A record passes if the Overall Score $\geq 0.85$, Anonymization Degree $\geq 0.90$, Content Consistency $\geq 0.90$, all dimensions $\geq 0.75$, and there are no privacy leaks in sensitive fields (name, address, email).
Table 4. Validation Thresholds Summary
Validation Stage Overall Score Critical Dimensions All Dimensions
Augmented Data $\geq$ 0.80
Logical Consistency $\geq$ 0.85
Data Completeness $\geq$ 0.80
Content Quality $\geq$ 0.80
$\geq$ 0.70
Privacy Rewriting $\geq$ 0.85
Anonymization Degree $\geq$ 0.90
Content Consistency $\geq$ 0.90
$\geq$ 0.75
3.1.4. Local Merchant Database Statistics
Overall, we constructed 1,354,185 representative samples from a pool of over 1,600,000 original seed records through data augmentation, privacy rewriting, and validation, which constitute our LocalSearchBench database.
The proportions of each category, the number of data for each city are presented in Figure 4(4(c))–4(4(k)). The distribution closely mirrors that of the raw data, ensuring our benchmark maintains the same categorical balance as real-world local life service platforms. The geographical distribution of merchant data for each city was derived through geospatial inference to generate latitude-longitude heatmaps, as shown in Figure 5. The distribution patterns align well with urban clustering phenomena and regional economic hierarchies, demonstrating the dataset’s realistic geographical representation.
(a) Standardized category distribution of real-world data
(b) City distribution
(c) Shanghai
(d) Beijing
(e) Guangzhou
(f) Shenzhen
(g) Hangzhou
(h) Chengdu
(i) Chongqing
(j) Wuhan
(k) Suzhou
Figure 4. City distribution and category distribution of 1,354,185 merchant data across 9 major cities in China.
(a) Shanghai
(b) Beijing
(c) Guangzhou
(d) Shenzhen
(e) Hangzhou
(f) Chengdu
(g) Chongqing
(h) Wuhan
(i) Suzhou
Figure 5. Geographical distribution heatmaps of 1,354,185 merchant data across 9 major cities in China. The color scale represents merchant density from 50 (blue) to 100 (red).
3.2. LocalRAG Construction
To support efficient and accurate merchant retrieval for LRMs, we build LocalRAG, a Retrieval-Augmented Generation system on our merchant database (Figure 8). LocalRAG uses a embedding model to convert each merchant profile into a high-dimensional vector. Structured fields are directly encoded into vectors and indexed in a city-partitioned vector database. During query processing, the target city is first extracted from the query text, then the corresponding city-partitioned vector database is selected. The query is embedded with the LocalRAG embedding model and matched via cosine similarity against merchant vectors from that city, selecting the top-N semantically related candidates through semantic retrieval. LocalRAG then applies a reranking model to the top-N candidates, scoring query-merchant relevance and returning the top-K merchants as the final retrieval results for LLM agents.
3.3. Multi-hop QA Construction
Figure 6. How Multi-hop QA pairs in LocalSearchBench are constructed.
In this subsection, we present the multi-hop QA construction of LocalSearchBench. Building upon the data construction pipeline described in the previous subsections, our multi-hop QA construction pipeline consists of four main stages: seed question collection, question instantiation, answer collection, and QA validation. Each of these stages is detailed in the following subsubsections.
3.3.1. Seed Question Collection
To ensure LocalSearchBench reflects real-world local life service complexities, we conducted a statistical analysis to understand user search behaviors and service interactions. Based on requirement complexity and reasoning demands, we classify questions into five intelligence levels (Table 5): L1 (Basic) for simple factual queries; L2 (Advanced) for ambiguous or multi-condition searches; L3 (Intelligent) for composite requirements with multi-turn reasoning; L4 (Expert) for personalized planning; L5 (AGI) for cross-platform coordination with adaptive reflection.
We collect 1,200 single-hop questions from local life service platform M (90%) and enriched sources (10%) across 6 service categories (in § 3.1.1). The platform data captures authentic user queries, while enriched data provides coverage for complex scenarios underrepresented in natural logs. Following the intelligence level taxonomy in Table 5, we focus on L3 (Intelligent) and L4 (Expert) questions, as they require multi-step reasoning within a static environment, while L5 (AGI) questions involve dynamic cross-platform interactions that are difficult to simulate offline.
(a) Intelligence Level Distribution
(b) Example Questions
Figure 7. Intelligence Level Analysis for Collected Questions
Table 5. Intelligence Level Classification for Seed Questions.
Intellig- ence Level Grade Requirement Understanding Planning-Search-Reflection Loop
L1 Basic
Standard expression of single
precise requirement
Supply matching search with $\leq$
2 conditions
L2 Advanced
Ambiguous expression of single
precise requirement;
Standard expression of
single ambiguous requirement
Supply matching search with $\geq$
3 conditions;
Semantic retrieval search supply;
Multi-source information
integration supply
L3 Intelligent
Composite requirements;
Multi-turn conversational
requirements
Reasoning with certain complexity;
Semantic retrieval with external
information dependency;
Procedural planning requirements;
Comparison and contrast scenarios
L4 Expert
Personalized requirements
External hard-to-access dependency;
Open-ended planning requirements
L5 AGI
Cross-platform coordination;
Complex decision-making;
Exception handling and reflection
Cross-platform dependency;
Task execution;
Adaptive reflection
3.3.2. Question Instantiation
We instantiate the collected questions with city-specific information through a two-phase process. First, we construct city profiles for each of the 9 cities (§ 3.1.1), collecting structured information from public resources including landmarks, cultural events, local cuisine, price bands, etc. To facilitate consistent evaluation, temporal elements are anchored to static timestamps, creating a fixed reference point for all queries. The collected city data is preprocessed using automated scripts to remove duplicates, normalize formats, and correct inconsistencies, then validated by annotators to ensure factual accuracy and cultural authenticity. Second, annotators from platform M manually instantiate questions into concrete multi-hop queries. For each city, approximately 100 questions are selected from the collected 1,200 questions and grounded with city-specific information. To transform single-hop questions into multi-hop queries, annotators enrich questions with 2–4 reasoning segments based on business relevance and logical coherence. These segments include: cross-merchant comparison, schedule planning, chained events, and constraint resolution. We maintain target intelligence-level and hop-count distributions (Figure 6), prioritizing underrepresented categories to ensure balanced coverage across service categories and cities in § 3.1.1.
3.3.3. Answer Collection
Figure 8. Overview of LocalPlayground.
For each instantiated question, we generate reference answers through a multi-step process using LocalRAG (§ 3.2), web search, and LLM-based generation. We first check whether the question needs real-time information (e.g., current events, weather, news) or can be answered with merchant data alone. If real-time signals are present, we invoke web search; otherwise, the query relies on the city-partitioned merchant vector database. For all questions, LocalRAG retrieves merchants scoped to the target city. For multi-hop questions, the hop count, retrieval target, and required search tool for each hop are predetermined during question construction. Search tools then execute iterative retrieval per hop, gathering merchants across categories/locations and combining them with any web-search snippets to form the full reasoning context. Claude-4.5-opus (Anthropic, 2025) then generate preliminary answers with expert oversight. The LLM reranks each hop’s retrieved candidates and selects the top-$k$ passages (k=5k{=}5) as evidence, then synthesizes a final answer grounded in these snippets while following the fixed hop plan and the original query. The prompt enforces: (1) grounding in retrieved evidence, (2) alignment with the fixed hop plan, (3) explicit citation of merchant and web evidence, and (4) avoidance of unsupported content. The resulting question-answer pairs serve as input for validation.
3.3.4. QA Validation
All preliminary QA pairs undergo a two-stage validation process (Figure 6). First, hop-by-hop validation ensures: (1) every hop has a clear answer, meaning retrieval targets must be found through LocalRAG or web search with sufficient evidence; (2) strict adherence to the target hop-count chain; (3) conformity to local life service business logic. Questions failing these criteria are manually modified and re-submitted to answer collection (§ 3.3.3). Second, three domain experts independently review answers for factual accuracy, completeness, and clarity, resolving disagreements through consensus. Answers are deterministic with balanced distribution across service categories and cities (§ 3.1.1), yielding 900 high-quality QA pairs for the final benchmark (Figure 6).
4. LocalPlayground
We design LocalPlayground for automated testing with LocalSearchBench. It contains a Search Agent for multi-hop retrieval and answering, and a Validation Agent for quality examination of answer and reasoning trajectory (Figure 8).
Search Agent. The Search Agent uses two tools: LocalRAG for vector-DB merchant retrieval and web search for real-time information (facts, weather, news, events) via Baidu APIs (Liu et al., 2021). It performs iterative search up to N rounds, calling each tool at most once per round to refine answers. For each query $q_{0}$, it records the trajectory $\tau=\{s_{0},a_{0},s_{1},a_{1},\dots,s_{t}\}$, with states $s_{i}=(q_{i},C_{i})$ storing the current sub-question and accumulated evidence. At step $i$, it selects a reasoning action $a_{i}^{\text{reason}}$ and, if needed, a search action $a_{i}^{\text{search}}$ to obtain evidence $E_{i+1}$, update $C_{i+1}=C_{i}\cup E_{i+1}$, and generate $q_{i+1}$ from $(q_{i},C_{i+1})$. A final reasoning step synthesizes $q_{t}$, and $s_{i+1}=F(s_{i},a_{i})$ forms a directed chain over sub-questions and evidence for reasoning.
Validation Agent. We employ an LLM-as-judge (Claude-4.5-opus) to evaluate each agent interaction along two dimensions: answer quality and trajectory quality. Answer quality is assessed on 7 metrics: correctness (factual accuracy), completeness (covers the query), fluency (readability), faithfulness (no hallucination), safety (no harmful content), avg. tool calls and avg. rounds. Each query is associated with a multi-attribute constraint checklist defining all necessary and sufficient conditions for a valid answer. During evaluation, the Validation Agent verifies whether the agent’s retrieved merchant satisfies every constraint on the checklist. If multiple merchants meet all listed conditions, each is considered a correct answer. Trajectory quality is scored on 4 dimensions: action relevance (step aligns with intent), evidence sufficiency (evidence is enough), causal coherence (later steps correctly reference earlier evidence), and efficiency (avoids redundant/ineffective calls).
Each instance is evaluated five times per dimension and averaged. Systems are anonymized, question order is randomized, and the judge model is distinct from all evaluated models.
Table 6. Performance comparison of different models on LocalSearchBench (Max Round N = 5). The top and worst performing results are highlighted in red (1st) and blue (bottom) backgrounds, respectively.
Models Answer Evaluation Metrics Trajectory Evaluation Metrics
Avg. tool calls Avg. rounds Correct- ness(%) Complete- ness(%) Fluency (%) Faithful- ness(%) Safety (%) Action Relevance (%) Evidence Sufficiency (%) Causal Coherence(%) Search Efficiency (%)
Qwen3-235B-A22B (w/o thinking) 2.00 2.93 21.20 50.94 69.16 25.28 79.72 75.22 43.61 50.68 45.99
Qwen3-235B-A22B (w/ thinking) 2.31 3.17 30.20 71.20 71.58 26.90 81.76 80.42 45.75 52.04 48.63
Qwen3-32B (w/o thinking) 2.78 3.11 19.80 40.96 68.50 21.38 80.76 74.67 46.52 48.82 49.84
Qwen3-32B (w/ thinking) 2.80 3.12 25.60 40.66 68.44 22.40 79.54 75.06 46.99 48.87 49.13
Qwen3-14B (w/o thinking) 2.53 2.07 24.20 40.60 69.62 27.44 80.78 80.46 46.44 50.96 50.02
Qwen3-14B (w/ thinking) 2.57 2.12 25.20 40.98 69.32 28.40 80.44 81.19 47.24 52.52 51.65
GPT-4.1 1.73 2.42 18.50 45.37 65.93 28.76 77.38 68.47 38.62 45.83 42.29
o3(high) 2.91 3.38 31.50 69.64 70.72 33.89 81.87 75.93 44.71 51.78 42.96
Gemini-2.5-Flash 1.84 2.51 21.00 58.44 68.04 35.72 79.70 70.58 41.61 48.94 46.91
Gemini-2.5-Pro 2.75 3.10 32.30 71.03 71.12 34.93 82.32 77.31 45.73 52.87 41.78
LongCat-Flash-Chat 2.34 3.07 25.30 52.98 69.45 27.49 83.61 77.78 47.33 50.86 52.29
LongCat-Flash-Thinking 3.04 3.20 30.70 68.83 69.07 31.47 80.10 78.50 47.37 53.18 53.27
GLM-4.6 (w/o thinking) 2.86 3.86 29.00 76.45 70.37 35.40 81.40 74.28 48.44 50.79 52.76
GLM-4.6 (w/ thinking) 3.08 4.06 32.80 76.83 70.27 37.48 81.30 77.66 48.90 52.67 54.43
Deepseek-V3.2 (w/o thinking) 3.19 4.18 31.70 77.08 70.12 35.96 81.52 75.51 48.49 52.23 54.33
Deepseek-V3.2 (w/ thinking) 3.21 4.20 35.60 77.56 70.92 39.78 81.13 75.58 48.86 52.62 54.83
Table 7. Ablation study of tool integration on thinking models using LocalSearchBench (Max Round N = 5). The top and worst performing results are highlighted in red (1st) and blue (bottom) backgrounds, respectively.
Methods Answer Evaluation Metrics Trajectory Evaluation Metrics
Avg. tool calls Avg. rounds Correct- ness(%) Complete- ness(%) Fluency (%) Faithful- ness(%) Safety (%) Action Relevance (%) Evidence Sufficiency (%) Causal Coherence(%) Search Efficiency (%)
LocalRAG workflow 1.00 1.00 15.18 58.30 70.80 66.10 85.40 37.70 23.53 32.77 39.35
Gemini-2.5-Pro (w/o tools) 0.00 1.00 0.00* 38.42 91.65 9.82 99.89 25.21 8.72 21.70 13.62
+ LocalRAG 1.86 2.83 23.11 63.92 82.32 38.65 84.95 68.10 37.37 46.93 39.96
+ LocalRAG & web search 2.75 3.10 32.30 71.03 71.12 34.93 82.32 77.31 45.73 52.87 41.78
GLM-4.6 (w/o tools) 0.00 1.00 0.00* 36.06 91.92 12.12 99.67 46.88 7.92 28.85 17.08
+ LocalRAG 2.33 3.32 23.91 69.21 83.75 41.45 84.72 63.33 36.21 48.02 52.05
+ LocalRAG & web search 3.08 4.06 32.80 76.83 70.27 37.48 81.30 77.66 48.90 52.67 54.43
Deepseek-V3.2 (w/o tools) 0.00 1.00 0.00* 39.30 90.20 18.56 99.55 52.65 10.80 32.10 41.15
+ LocalRAG 2.88 3.87 27.04 70.05 82.98 40.59 83.85 65.55 42.78 47.85 54.07
+ LocalRAG & web search 3.21 4.20 35.60 77.56 70.92 39.78 81.13 75.58 48.86 52.62 54.83
* The retrieved merchant must have a poi_id from LocalRAG.
Table 8. Sensitivity analysis of Max Round N on model performance. The top and worst performing results are highlighted in red (1st) and blue (bottom) backgrounds, respectively.
Max Round Answer Evaluation Metrics Trajectory Evaluation Metrics
Avg. tool calls Avg. rounds Correct- ness(%) Complete- ness(%) Fluency (%) Faithful- ness(%) Safety (%) Action Relevance (%) Evidence Sufficiency (%) Causal Coherence(%) Search Efficiency (%)
$N=2$ 2.46 2.82 30.50 70.61 76.44 54.65 89.94 59.94 31.17 44.82 40.77
$N=3$ 2.78 3.30 33.50 74.79 73.57 43.50 84.58 61.68 42.96 47.42 48.50
$N=4$ 3.03 3.95 34.20 76.34 72.27 41.26 82.53 72.97 49.28 51.32 54.13
$N=5$ 3.21 4.20 35.60 77.56 70.92 39.78 81.13 75.58 48.86 52.62 54.83
$N=6$ 3.76 4.75 34.90 74.52 70.36 38.56 80.91 68.63 49.28 51.65 47.10
$N=7$ 3.87 4.85 33.70 73.60 70.19 37.36 80.84 69.50 49.37 48.20 44.99
$N=8$ 4.02 5.01 32.40 73.15 70.24 37.35 80.39 66.89 49.39 47.94 43.58
5. Evaluation
5.1. Evaluated Models
We evaluate 16 state-of-the-art thinking and non-thinking models for agents: GPT series (GPT-4.1, o3 (high)) (OpenAI, 2025), Gemini series (Gemini-2.5-Flash, Gemini-2.5-Pro) (Comanici et al., 2025), LongCat series (LongCat-Flash-Chat, LongCat-Flash-Thinking) (Team et al., 2025), Deepseek-V3.2 (DeepSeek-AI, 2025), GLM-4.6 (Zeng et al., 2025), and Qwen3 series (Qwen3-14B, Qwen3-32B, Qwen3-235B-A22B) (Yang et al., 2025a). Unless specified, LocalPlayground uses default settings: maximum tool call rounds N = 5; LocalRAG with Qwen3-Embedding-8B and Qwen3-Reranker-8B (Zhang et al., 2025c) (top-100 dense retrieval, top-20 reranking); web search top-k = 20; and LLM temperature = 0. Experiments run on Intel Xeon Gold 5218 @ 2.30GHz with 1 $\times$ H20-141G GPU.
5.2. Data and LocalRAG Quality
We evaluate LocalSearchBench database with the two-stage framework in §3.1.3, sampling 1,000 merchants. As shown in Table 18, the augmented and privacy-rewritten data achieve high overall quality scores of 0.86 and 0.92 (out of 1).
[…]
Since retrieval quality directly affects downstream QA performance, we evaluate LocalRAG on 27,360 hop-level queries using NDCG@10/20 and MRR@10/20 over the top-20 results of the full retrieval pipeline (§3.2). Our best configuration, Qwen3-Embedding-8B, reaches NDCG@10/20 of 0.84/0.87 and MRR@10/20 of 0.82/0.83 (Table 9). These hop-level queries come from the multi-hop QA construction in §3.3 with ground-truth evidence for each hop; metric definitions and full retrieval results are detailed in Appendix A.1.
5.3. Overall Performance
The choice of underlying LRM significantly influences performance across all metrics. Deepseek-V3.2 achieves the highest Correctness (35.60%), followed by GLM-4.6 (32.80%) and Gemini-2.5-Pro (32.36%). It also leads in Completeness (77.56%) and Faithfulness (39.78%). Thinking models outperform non-thinking ones in answer quality metrics (Correctness: 30.49% vs. 21.34%, Completeness: 64.59% vs. 49.10%) and most trajectory evaluation metrics , with similar Search Efficiency. Despite these variations, overall performance remains limited, with the best model achieving only 35.60% Correctness, underscoring the substantial challenge of multi-hop local search. Across all 16 models, correctness further drops on L4 tasks (23.51%) compared to L3 (28.66%), reflecting the greater challenge of L4 (see Appendix A.6).
Human evaluation. While LLM-as-judge provides scalable assessment, it may introduce biases. Our human evaluation validates its reliability, showing high consistency with human judgment in relative model ranking. We conduct two studies: (1) 3 annotators evaluate 900 outputs from Deepseek-V3.2, and (2) 100 randomly sampled outputs each from GLM-4.6, Gemini-2.5-Pro, LongCat-Flash-Thinking, and Qwen3-235B-A22B. Human scores demonstrate substantial inter-rater reliability (Cohen’s $\kappa$ = 0.79, 95% CI: $[0.75,0.83]$, 87.91% raw agreement). Detailed results are in Appendix A.2.
5.4. Ablation Study
We conduct ablation study on tool integration–a core design element for agentic search in local life services. Specifically, we evaluate three configurations across top-performing LRMs (Gemini-2.5-Pro, GLM-4.6, Deepseek-V3.2): (1) LRM-only (no tools), (2) LRM + LocalRAG (merchant retrieval only), and (3) LRM + LocalRAG + web search (full toolchain). The goal is to quantify how each tool impacts answer and trajectory quality.
As shown in Table 7, incorporating web search into LocalPlayground improves answer quality but reduces faithfulness. On average, it boosts Correctness by 8.90 pp and Completeness by 7.41 pp, while decreasing Faithfulness by 2.83 pp. This trade-off is most evident in Gemini-2.5-Pro, which gains 9.25 pp in Correctness but loses 3.72 pp in Faithfulness. These results highlight web search’s dual role in enabling critical information retrieval and introducing noise in multi-hop reasoning.
5.5. Sensitivity Study
To investigate the impact of maximum conversation rounds $N$ on tool usage and model performance, we conduct a sensitivity study using Deepseek-V3.2, the best-performing model, with varying $N$ from 2 to 8. As shown in Table 8, the optimal configuration is $N=5$, achieving the highest Correctness (35.60%). When $N<5$, insufficient rounds limit information gathering, resulting in lower correctness (30.50%–34.20%). When $N>5$, despite increased tool usage, correctness declines (34.90%–32.48%), suggesting excessive rounds introduce noise that hinders answer quality. In terms of computational cost, tool calls and conversation length increase with $N$, peaking at $N=8$ (4.02 tool calls, 5.01 rounds). Therefore, $N=5$ represents the optimal balance of performance and efficiency.
5.6. In-Depth Analysis
Figure 9. Failure Case Studies. A typical chain event: the 2nd and 3rd hops depend on the 1st hop.
We study factors affecting LRM performance on agentic search tasks in local life services. Figure 9 shows concrete examples.
Unstable tool call strategy (38.6%). Lack of a unified strategy for tool call timing and frequency leads to insufficient tool calls or wasted budget. LRMs may also violate tool-use rules, such as using web search for merchant queries or making redundant tool calls within a single dialogue turn.
Missing explicit multi-hop planning (30.9%). LRMs lacking hop sequences fail to decompose tasks for retrieval or tool use. Without explicit hop targets, sub-goals remain unordered, causing empty matches and higher coverage failures in multi-hop tasks. Retrieval failures cascade through steps due to the absence of validation or backtracking.
Query generation dilutes critical constraints (11.8%). Naively generated long-form queries often omit key constraints, weakening retrieval signals and reducing recall of relevant candidates.
Long-context noise (18.7%). Lengthy search results and information sent to LRMs bring irrelevant data, reduce the ability to distinguish relevant results, and can exceed context window limits.
6. Conclusion
We propose LocalSearchBench, a benchmark for agentic search in local life services evaluating multi-hop reasoning. The benchmark contains 1,354,185 database and 900 multi-hop QA tasks across 9 Chinese cities and 6 service categories. Experiments on 16 leading LRMs using LocalPlayground show current models struggle, with DeepSeek-V3.2 achieving only 35.60% correctness. We open-source LocalSearchBench and LocalPlayground to advance domain-specific agentic search.
Limitations
We highlight two main limitations.
Dataset construction and agent workflows. Our dataset construction and agent workflows rely on large, often closed-source LLMs and human-in-the-loop verification. This dependence improves generation quality but raises reproducibility and cost concerns, consuming approximately 200M tokens and costing approximately \$1,200 for API usage, plus over 220 hours of human labor for verification; open-source substitutes require substantially more compute and operational effort to match performance.
Language and Locale Specificity. The benchmark is monolingual (Chinese) and geographically scoped to mainland China. This design prioritizes domain depth, authenticity, and precise geospatial grounding, which are critical for evaluating complex, real-world planning. However, it inherently limits direct applicability to other languages and regions, posing a challenge for models without relevant language or domain knowledge. The core framework LocalPlayground, including its task formulation, evaluation suite, and tool-integrated environment, remains fully generalizable. Therefore, this work provides a necessary, localized instantiation and a clear methodological template for the community to build and compare culturally adapted benchmarks.
Appendix A Evaluation Results and Performance Analysis
A.1. Ranking Metrics Definitions
For completeness, we summarize the ranking metrics used to evaluate retrieval quality. Given a set of hop-level queries $Q$ and a cutoff $k\in\{10,20\}$, NDCG@$k$ and MRR@$k$ are defined as
(1) $\small\mathrm{NDCG@}k=\frac{1}{|Q|}\sum_{q\in Q}\frac{\mathrm{DCG@}k(q)}{\mathrm{IDCG@}k(q)},$
with
(2) $\small\mathrm{DCG@}k(q)=\sum_{i=1}^{k}\frac{2^{\mathrm{rel}_{q,i}}-1}{\log_{2}(i+1)},$
and
(3) $\small\mathrm{MRR@}k=\frac{1}{|Q|}\sum_{q\in Q}\frac{1}{\mathrm{rank}_{q}},$
where $\mathrm{rel}_{q,i}$ is the relevance label of the $i$-th retrieved document for query $q$, and $\mathrm{rank}_{q}$ is the rank of the first relevant document within the top-$k$ results (0 if none is found).
For hop-level retrieval evaluation, relevance labels are assigned via rule-based automatic annotation with human sampling validation. A merchant is labeled as relevant ($\mathrm{rel}_{q,i}$=1) if it fully satisfies all constraints of the hop query (e.g., geographic scope, service category, rating threshold); otherwise, it is labeled as irrelevant ($\mathrm{rel}_{q,i}$=0). All relevant merchants constitute the ground-truth evidence set for the query–retrieval is considered effectively recalled if any ground-truth merchant is included in the results, and ranking quality is measured by NDCG@k to reflect the position of ground-truth merchants. To ensure reliability, 10% of labels are manually validated by two domain experts, with inter-annotator agreement Cohen’s $\kappa$ = 0.81 (95% CI: [0.78, 0.84]).
We evaluate LocalRAG’s retrieval stage (retrieval + reranker) on 27,360 hop-level queries. Embeddings are produced by the respective embedding models listed below, while the reranker is fixed to Qwen3-Reranker-8B for all configurations. The metrics reported are NDCG@10/20 and MRR@10/20 computed over the top-20 candidates returned by the full pipeline (dense retrieval followed by reranking). Table 9 summarizes overall retrieval quality across embedding backbones.
Table 9. Overall retrieval quality of LocalRAG on hop-level queries.
Method NDCG@10 NDCG@20 MRR@10 MRR@20
BM25 0.60 0.64 0.55 0.57
Qwen3-Embedding-8B 0.84 0.87 0.82 0.83
Qwen3-Embedding-4B 0.78 0.81 0.75 0.76
Qwen3-Embedding-0.6B 0.70 0.74 0.66 0.68
BGE-M3 0.72 0.75 0.67 0.68
Independent human relevance validation
We note that rule-based labels are generated within the same pipeline that instantiates hop queries and collects evidence, which can create coupling between ground-truth signals and the retrieval system. To quantify this effect, we performed an independent human validation on a stratified sample of 1,000 hop-level queries (stratified by city, hop-count and service category). Three domain experts annotated the top-20 retrievals per query (blind to pipeline labels); majority vote produced the human reference labels.
For the sampled subset we compute NDCG@10/20 and MRR@10/20 under both the pipeline rule-based labels and the independent human labels, report inter-annotator Cohen’s $\kappa$, and measure Spearman rank correlation of model rankings between the two label sets. Table 10 is a template for reporting these results.
Table 10. Comparison of retrieval metrics between independent human judgments and pipeline-derived relevance labels
Metric Rule labels Human labels Delta (abs) Delta (%)
NDCG@10 0.842 0.838 0.004 0.48%
NDCG@20 0.871 0.868 0.003 0.34%
MRR@10 0.821 0.818 0.003 0.37%
MRR@20 0.832 0.829 0.003 0.36%
Inter-annotator Cohen’s $\kappa$ 0.81 (95% CI: [0.78, 0.84])
Spearman rank $\rho$ (rule vs human) 0.93
If the observed deltas are small (e.g., absolute $\Delta<1\%$ and Spearman $\rho\geq 0.90$), we treat the pipeline-derived labels as robust for retrieval evaluation; larger discrepancies are reported and discussed in the main text.
A.2. Human Evaluation Results
To complement the automated LLM-as-judge evaluation, we conduct human evaluation on the best-performing model from our experiments. Three experienced annotators independently assess the outputs of Deepseek-V3.2 (w/ thinking) on a sampled subset of LocalSearchBench, evaluating both answer quality and trajectory quality using the same metrics as our automated evaluation framework. The detailed human evaluation results are presented in Table 11.
Table 11. Human evaluation results of Deepseek-V3.2 on LocalSearchBench (Max Round N = 5).
Model Answer Evaluation Metrics Trajectory Evaluation Metrics
Avg. tool calls Avg. rounds Correct- ness(%) Complete- ness(%) Fluency (%) Faithful- ness(%) Safety (%) Action Relevance (%) Evidence Sufficiency (%) Causal Coherence(%) Search Efficiency (%)
Deepseek-V3.2 (w/ thinking) 3.21 4.20 33.50 75.42 69.18 36.94 79.67 72.85 47.29 51.06 53.71
To quantify the agreement between human evaluators, we compute weighted Cohen’s $\kappa$ for each metric across the three annotators. The average $\kappa$ across all metrics is 0.79 (95% CI: [0.75, 0.83]), corresponding to 87.91% raw agreement, indicating substantial inter-rater reliability.
The weighted Cohen’s kappa is computed as
(4) $\kappa_{w}=\frac{P_{o}-P_{e}}{1-P_{e}},$
where
(5) $P_{o}=\sum_{i,j}w_{ij}p_{ij},\qquad P_{e}=\sum_{i,j}w_{ij}p_{i\cdot}p_{\cdot j},$
with $p_{ij}$ the observed proportion of ratings in category $i$ by rater A and $j$ by rater B, $p_{i\cdot}$ and $p_{\cdot j}$ the marginal proportions, and $w_{ij}$ the weights. We use quadratic weights
(6) $w_{ij}=1-\frac{(i-j)^{2}}{(K-1)^{2}},$
for $K$ ordered categories. For multiple annotators we average pairwise weighted kappas.
A.3. Retrieval and Reranking Sensitivity Analysis
To investigate how sensitive our evaluation results are to the choice of retriever and reranker components, we conduct a comprehensive sensitivity analysis across multiple retrieval configurations. This analysis is crucial for understanding the robustness of our findings and provides guidance for practitioners selecting retrieval systems for agentic search tasks.
Experimental Setup: We evaluate the performance of DeepSeek-V3.2 (the best-performing model in our primary evaluation) across eight different retrieval configurations, combining four embedding models with two reranking strategies:
  • Embedding Models: Qwen3-Embedding-8B, bge-m3, Qwen3-Embedding-4B, Qwen3-Embedding-0.6B
  • Reranking Strategies: Qwen3-Reranker-8B (our primary reranker), and a baseline no-reranking approach using only embedding similarity scores
Each configuration uses identical search parameters (top-20 retrieval with reranking when applicable) and is evaluated on the same 900 test instances from our benchmark.
Performance Variation Across Configurations: Table 12 summarizes the performance of DeepSeek-V3.2 across different retrieval configurations, measured by key answer-level and trajectory-level metrics.
Table 12. DeepSeek-V3.2 performance across different retrieval configurations
Configuration Correctness Completeness Faithfulness Evidence Causal Search
(%) (%) (%) Sufficiency (%) Coherence (%) Efficiency (%)
Qwen3-8B + Qwen3-Reranker 35.60 77.56 39.78 48.86 52.62 54.83
bge-m3 + Qwen3-Reranker 32.85 74.23 36.91 45.12 49.78 51.94
Qwen3-4B + Qwen3-Reranker 31.42 72.89 35.67 43.85 48.92 50.23
Qwen3-0.6B + Qwen3-Reranker 28.76 69.45 32.34 41.23 46.78 47.56
Qwen3-8B + No Reranking 31.23 73.89 35.42 44.67 48.91 57.12
bge-m3 + No Reranking 28.94 71.34 33.18 42.45 47.23 54.67
Qwen3-4B + No Reranking 27.51 68.92 31.89 40.78 45.67 52.89
Qwen3-0.6B + No Reranking 25.34 65.78 29.45 38.92 43.56 50.34
Key Findings on Retrieval Sensitivity:
  1. (1)
    Embedding Model Impact: Performance varies significantly across embedding backbones, with Qwen3-Embedding-8B achieving 35.60% correctness compared to 28.76% for Qwen3-Embedding-0.6B–a relative difference of 23.8%. This suggests that embedding quality directly influences downstream agentic reasoning capabilities.
  2. (2)
    Reranking Benefits: The Qwen3-Reranker-8B consistently improves performance across all metrics and embedding models. For instance, reranking boosts correctness by 4.37 percentage points on average (from 28.26% to 32.63% without reranking). However, reranking comes at a computational cost, reducing Search Efficiency by approximately 4-5 percentage points.
  3. (3)
    Trade-offs and Optimal Configurations: The optimal configuration (Qwen3-8B + Qwen3-Reranker) achieves the highest correctness and evidence quality but at the cost of reduced search efficiency. Configurations without reranking show higher efficiency but lower answer quality, suggesting a clear trade-off between computational cost and performance.
  4. (4)
    Metric-Specific Sensitivity: Different metrics show varying sensitivity to retrieval quality. Correctness and Faithfulness are most sensitive (variation range: 10.26% and 10.33% respectively), while Search Efficiency shows relatively stable performance across configurations.
Implications for Robustness Analysis: Our sensitivity analysis reveals that evaluation results can vary by up to 24% depending on retrieval configuration choice. This underscores the need for systematic robustness studies that evaluate agentic search systems across multiple retrieval backbones rather than relying on single configurations.
This sensitivity analysis demonstrates that while our primary findings hold across multiple configurations, the choice of retrieval system significantly impacts quantitative results. Practitioners should consider these trade-offs when deploying agentic search systems in real-world applications.
A.4. Judge Model Robustness Analysis
To address potential bias from using the same model family (Claude-4.5-opus) for both ground truth generation and evaluation, we conduct a comprehensive robustness analysis using two independent judge models: Claude-4.5-opus and GPT-4-turbo. This analysis evaluates whether model rankings remain stable across different judges, quantifies inter-judge agreement, and assesses the calibration of absolute scores.
Evaluation Setup: We evaluate the same set of 900 test instances across all models using both judge models. Each judge independently scores all metrics using identical prompts and evaluation criteria. This comprehensive analysis covers all 14 model configurations from our primary evaluation, enabling us to assess both ranking stability and absolute score calibration across the full model spectrum.
Inter-Judge Agreement: The correlation between Claude-4.5-opus and GPT-4-turbo scores across all metrics and models is 0.89 (Spearman’s $\rho$, $p<0.001$), indicating strong agreement. Agreement varies across metrics, with highest consistency for Correctness ($\rho=0.94$) and relatively lower for Faithfulness ($\rho=0.82$). The high agreement level suggests that both judges apply evaluation criteria consistently, despite potential differences in their training data and architectural approaches.
Comprehensive Model Performance Across Judges: Table 13 presents the performance comparison across judge models for all evaluated models, ranked by overall performance under Claude-4.5-opus.
Table 13. Comprehensive model performance comparison across judge models
Model Claude-4.5-opus GPT-4-turbo Rank
Correctness Completeness Overall Correctness Completeness Overall Stability
DeepSeek-V3.2 (w/ thinking) 35.60 77.56 56.58 34.85 76.92 55.89
GLM-4.6 (w/ thinking) 32.80 76.83 54.82 32.12 76.45 54.29
Gemini-2.5-Pro (w/ thinking) 32.30 71.03 51.67 31.89 70.78 51.34
Qwen3-235B-A22B (w/ thinking) 30.20 71.20 50.70 29.87 70.91 50.39
LongCat-Flash-Thinking 30.70 68.83 49.77 30.45 68.67 49.56
Gemini-2.5-Flash 21.00 58.44 39.72 20.67 58.12 39.40
Qwen3-32B (w/ thinking) 25.60 40.66 33.13 25.34 40.45 32.90
DeepSeek-V3.2 (w/o thinking) 31.70 77.08 54.39 31.23 76.78 54.01
GLM-4.6 (w/o thinking) 29.00 76.45 52.73 28.67 76.12 52.40
LongCat-Flash-Chat 25.30 52.98 39.14 25.01 52.67 38.84
Qwen3-235B-A22B (w/o thinking) 21.20 50.94 36.07 20.89 50.67 35.78
Qwen3-14B (w/ thinking) 25.20 40.98 33.09 24.89 40.73 32.81
Qwen3-32B (w/o thinking) 19.80 40.96 30.38 19.56 40.78 30.17
Qwen3-14B (w/o thinking) 24.20 40.60 32.40 23.94 40.35 32.15
Model Ranking Stability: The ranking stability across judges is remarkably high. The top-5 model rankings remain identical across both judges, with only minor reordering within the middle ranks. Key observations include:
  1. (1)
    Top Performer Consistency: DeepSeek-V3.2 (with thinking) maintains its top position across both judges, demonstrating robust superiority in agentic search capabilities.
  2. (2)
    Thinking vs. Non-Thinking Gap: The performance gap between thinking and non-thinking variants is consistently observed across judges, with thinking models showing 3-5% higher correctness scores.
  3. (3)
    Family-Specific Patterns: Models within the same family (e.g., Qwen3 variants, Gemini variants) maintain their relative ordering, suggesting that judge models agree on family-level capabilities.
Absolute Score Calibration: Beyond ranking stability, we analyze the calibration of absolute scores between judges. The mean absolute difference in correctness scores is 0.42% (SD = 0.31%), indicating excellent calibration. For completeness scores, the mean difference is 0.35% (SD = 0.28%). These small differences suggest that both judges apply similar scoring standards, with no systematic bias toward over- or under-estimation.
Metric-Specific Agreement Analysis: Table 14 breaks down inter-judge agreement by metric type, revealing nuanced patterns in evaluation consistency.
Table 14. Inter-judge agreement across all evaluation metrics
Metric Spearman’s $\rho$ Mean Abs. Diff. Max Abs. Diff.
Correctness 0.94 0.42% 0.75%
Completeness 0.91 0.35% 0.61%
Faithfulness 0.82 1.45% 2.12%
Fluency 0.87 0.53% 0.89%
Safety 0.95 0.26% 0.42%
Overall Agreement 0.89 0.60% 2.12%
Implications for Benchmark Reliability: The comprehensive judge robustness analysis validates the reliability of our evaluation framework across multiple dimensions:
  1. (1)
    Ranking Stability: Model rankings are highly stable across judges, providing confidence in comparative evaluations.
  2. (2)
    Absolute Score Calibration: The small absolute differences indicate that scores are well-calibrated and not subject to systematic judge-specific biases.
  3. (3)
    Metric Consistency: All evaluation metrics demonstrate strong inter-judge agreement, with varying degrees of consistency across different aspects of agentic search performance.
  4. (4)
    Framework Validation: The use of Claude-4.5-opus for both ground truth generation and evaluation does not introduce systematic bias that affects model comparisons.
This extended analysis demonstrates that our evaluation results are robust not only for the best-performing model but across the entire spectrum of evaluated systems, providing strong validation of the benchmark’s reliability and discriminatory power.
A.5. Comprehensive Retrieval Performance Across Data Processing Stages
To fully validate the data augmentation and privacy rewriting pipeline, we conduct a comprehensive experiment comparing retrieval performance across three data processing stages: (1) original seed data, (2) augmented data, and (3) privacy-rewritten data.
Experimental Setup: Using the same 10,000 merchant records and 1,000 search queries as in the privacy validation experiment, we evaluate retrieval performance at each processing stage using identical retrieval configurations (Qwen3-Embedding-8B + Qwen3-Reranker-8B).
Data Processing Stages:
  • Original Seed Data: Raw merchant records with only 12 basic fields (name, category, address, etc.)
  • Augmented Data: Enhanced records with 29 fields through generative augmentation
  • Privacy-Rewritten Data: Final dataset with sensitive fields anonymized while preserving business information
Results: The augmentation process significantly improves retrieval quality by enriching merchant profiles with detailed attributes, while privacy rewriting maintains this enhanced performance with minimal degradation.
Table 15. Retrieval performance across data processing stages
Metric Original Seed Augmented Privacy-Rewritten
NDCG@10 0.723 0.842 0.839
NDCG@20 0.752 0.871 0.868
MRR@10 0.684 0.821 0.818
MRR@20 0.698 0.832 0.829
Performance Gains: Data augmentation improves NDCG@10 by 16.5% (from 0.723 to 0.842), demonstrating that enriched merchant profiles provide substantially better retrieval signals. Privacy rewriting causes only 0.36% average degradation, preserving 98.9% of the augmented performance.
Implications: This comprehensive evaluation confirms that our data processing pipeline successfully enhances retrieval quality through augmentation while maintaining performance through careful anonymization, ensuring the benchmark supports realistic and effective search scenarios.
A.6. Full Model Performance Table
For completeness, we provide the full per-metric performance comparison of all models, including L3 and L4 task breakdowns, in Table 16.
Overall best performer: Deepseek-V3.2 (w/ thinking) achieves the strongest overall profile on our automated metrics, registering the highest average Correctness (35.60%), the top Completeness score (77.56%), and the highest Search Efficiency (54.83%). These results indicate that the model is both effective at satisfying checklist constraints and relatively efficient at locating supporting evidence during the multi-hop process.
Thinking vs. non-thinking behavior: Enabling the chain-of-thought / “thinking” setting generally improves correctness and completeness for most models, often by a substantial margin (several percentage points). This improvement comes at the cost of slightly increased interaction cost: thinking variants typically use more tool calls and more rounds on average. The trade-off suggests that controlled deliberation helps planning and evidence aggregation, but requires careful engineering to avoid excessive search overhead.
Difficulty breakdown (L3 vs. L4): Performance is consistently higher on L3 tasks than on L4 tasks across models and metrics. For example, Deepseek-V3.2 (w/ thinking) shows a notable Correctness gap between L3 (38.70%) and L4 (28.90%). This confirms that L4 instances–characterized by personalized, hard-to-access dependencies and open-ended planning–pose substantially greater challenges for both retrieval and reasoning.
Trajectories and evidence quality: Trajectory-level metrics reveal complementary strengths and weaknesses. Some systems (e.g., LongCat variants) achieve very high Action Relevance and Causal Coherence, indicating strong procedural planning, while others (including the top-ranked Deepseek) balance high Completeness and Faithfulness with strong Search Efficiency. Across the board, Evidence Sufficiency remains moderate (typically in the mid-40s to low-50s %), showing that even high-scoring models often fail to gather fully sufficient, verifiable supporting evidence for every checklist item.
Metric correlations and practical implications: Higher Completeness and Faithfulness tend to correlate with more rounds and tool calls, suggesting a trade-off between answer thoroughness and interaction cost. The L4 difficulty gap and the moderate Evidence Sufficiency scores point to two priority directions for improvement: (1) stronger, hop-aware retrieval/reranking to surface more task-relevant evidence early, and (2) improved multi-hop planning and tool-selection heuristics to avoid wasted queries and to better synthesize retrieved passages into verifiable checklist items.
These observations synthesize the per-model trade-offs visible in Table 16 and provide actionable guidance for improving agentic search pipelines on challenging, real-world multi-hop tasks.
Checklist derivation, normalization and scoring
To improve reproducibility and interpretability of the answer-level metrics (Correctness, Completeness, Faithfulness) reported in Table 16, we summarize how the multi-attribute constraint checklist is derived, normalized across tasks, and used for cross-task scoring.
Checklist derivation. For each instantiated multi-hop task we extract explicit hard constraints (e.g., city, district, service category, minimum rating, time window) and optional preferences (e.g., price preference). The checklist is composed of:
  • Hard constraints: conditions that must be strictly satisfied for the instance to be counted as correct (e.g., “landmark = Xujiahui”, “rating $\geq$ 3.5”).
  • Soft constraints: preferences that affect completeness or quality but do not by themselves invalidate correctness (e.g., “price range = mid”).
  • Evidence items: minimal set of verifiable evidence pieces required to validate each checklist entry (see Evidence requirements).
Field normalization. To ensure consistent comparison across tasks and sources we normalize merchant fields before checking constraints:
  • Business hours: normalized to 24h intervals (e.g., 08:00–22:00) and parsed to start/end timestamps for interval checks.
  • Price fields: mapped to RMB numeric ranges or coarse levels (low/medium/high) with task templates defining comparison operators.
  • Address / administrative region: standardized to a (city, district, business_area, landmark) quadruplet;
Conflict and missing-field handling. When automated checks encounter conflicts or missing values we apply deterministic rules:
  1. (1)
    If a hard constraint is violated or cannot be verified, the instance is marked as failing Correctness and flagged for manual review (used to calibrate annotators).
  2. (2)
    If a soft constraint is partially met, it receives partial credit (see scoring) rather than invalidating correctness.
  3. (3)
    If a field is missing but can be inferred reliably from other fields (e.g., city inferred from address), we use the inferred value and record the inference source; otherwise the item is treated as “insufficient evidence”.
Evidence requirements. Each checklist item must be supported by at least one verifiable evidence item drawn from the recorded retrieval trajectory (LocalRAG/Web search). Accepted evidence types:
  • Merchant record fields with a stable identifier (poi_id).
  • Extracted web snippets with timestamps.
  • Retrieved RAG passages that explicitly contain the required attribute.
The minimal evidence set for a positive verification typically includes the merchant identifier plus the specific field value and its source (e.g., rating=4.2 from LocalRAG: poi_bj_12345678).
Scoring and aggregation. We evaluate attributes with simple, transparent rules:
  • Attributes are scored per-attribute as pass/fail (1/0).
  • Dimension-level scores are normalized to percentages (0–100%) and combined across tasks using weighted averages (task weights reflect task difficulty and importance).
  • Correctness is binary at the instance level: all key hard constraints must be satisfied to assign Correctness = 100%, otherwise Correctness = 0%.
Annotator calibration and quality control. To guarantee cross-annotator consistency:
  • Annotators receive a detailed handbook with rules, examples, and exception handling.
  • Each annotator must pass a calibration test on 100 gold-standard instances (pass threshold $\geq$ 90%) before labeling.
  • Disagreements are resolved by majority vote or expert arbitration; we periodically re-check 5% of samples to detect drift.
Illustrative checklist examples. Representative task-level checklists and the expected evidence/answer format are provided to annotators (see Appendix A.5 for sample formats). Example templates include required fields, evidence snippets, and rationale fields to improve auditability and reproducibility.
Table 16. Performance comparison of different models on LocalSearchBench (Max Round N = 5) with L3 and L4 task breakdowns. The top and worst performing results are highlighted in red (1st) and blue (bottom) backgrounds, respectively.
Models Answer Evaluation Metrics Trajectory Evaluation Metrics
Avg. tool calls Avg. rounds Correct- ness(%) Complete- ness(%) Fluency (%) Faithful- ness(%) Safety (%) Action Relevance (%) Evidence Sufficiency (%) Causal Coherence(%) Search Efficiency (%)
Avg. L3 L4 Avg. L3 L4 Avg. L3 L4 Avg. L3 L4 Avg. L3 L4 Avg. L3 L4 Avg. L3 L4 Avg. L3 L4 Avg. L3 L4 Avg. L3 L4 Avg. L3 L4
Qwen3-235B-A22B (w/o thinking) 2.00 2.04 1.92 2.93 2.97 2.84 21.20 24.90 13.20 50.94 51.14 50.54 69.16 69.53 68.42 25.28 27.57 20.70 79.72 80.00 79.16 75.22 77.57 70.52 43.61 44.43 41.97 50.68 52.51 47.02 45.99 47.04 43.89
Qwen3-235B-A22B (w/ thinking) 2.31 2.40 2.13 3.17 3.21 3.10 30.20 31.10 28.30 71.20 72.29 69.02 71.58 71.85 71.04 26.90 28.62 23.46 81.76 82.08 81.12 80.42 83.93 73.40 45.75 46.91 43.43 52.04 53.11 49.90 48.63 50.67 44.55
Qwen3-32B (w/o thinking) 2.78 2.82 2.70 3.11 3.09 3.16 19.80 22.00 15.10 40.96 40.98 40.92 68.50 68.81 67.88 21.38 22.90 18.34 80.76 80.85 80.58 74.67 80.41 63.19 46.52 47.79 43.98 48.82 50.71 45.04 49.84 53.28 42.96
Qwen3-32B (w/ thinking) 2.80 2.87 2.65 3.12 3.09 3.18 25.60 25.80 25.20 40.66 40.53 40.92 68.44 68.83 67.66 22.40 24.28 18.64 79.54 79.30 80.02 75.06 80.04 65.10 46.99 48.66 43.65 48.87 51.08 44.45 49.13 51.52 44.35
Qwen3-14B (w/o thinking) 2.53 2.64 2.30 2.07 2.02 2.17 24.20 25.20 22.20 40.60 41.26 39.28 69.62 69.65 69.56 27.44 29.59 23.14 80.74 80.47 81.28 80.46 85.50 70.38 46.44 48.00 43.32 50.96 53.20 46.48 50.02 53.28 43.50
Qwen3-14B (w/ thinking) 2.57 2.70 2.31 2.12 2.07 2.21 25.20 27.00 21.60 40.98 41.20 40.54 69.32 69.79 68.38 28.40 31.29 22.62 80.44 80.70 79.92 81.19 84.88 73.81 47.24 48.16 45.40 52.52 53.01 51.54 51.65 53.36 48.23
GPT-4.1 1.73 1.86 1.47 2.42 2.58 2.11 18.50 19.80 15.90 45.37 46.82 42.47 65.93 66.24 65.32 28.76 30.41 25.37 77.38 77.69 76.76 68.47 71.28 62.94 38.62 39.73 36.40 45.83 47.26 43.17 42.29 43.91 39.45
o3(high) 2.91 3.04 2.65 3.38 3.47 3.21 31.50 32.90 28.80 69.64 70.38 68.26 70.72 71.03 70.11 33.89 35.72 30.42 81.87 82.19 81.23 76.93 78.67 73.85 44.71 45.83 42.47 51.78 52.94 49.46 42.96 44.18 40.52
Gemini-2.5-Flash 1.84 1.99 1.52 2.51 2.65 2.22 21.00 23.80 15.40 58.44 59.71 55.90 68.04 67.68 68.76 35.72 39.85 27.46 79.70 80.06 78.98 70.58 73.86 64.02 41.61 42.86 39.11 48.94 50.09 46.64 46.91 49.24 42.25
Gemini-2.5-Pro 2.75 2.81 2.64 3.10 3.17 2.95 32.30 33.80 29.30 71.03 71.42 70.25 71.12 71.20 70.96 34.93 37.74 29.31 82.32 82.26 82.44 77.31 78.49 74.95 45.73 46.06 45.07 52.87 54.81 48.99 41.78 43.00 39.34
LongCat-Flash-Chat 2.34 2.36 2.30 3.07 3.03 3.15 25.30 25.60 24.60 52.98 55.76 47.42 69.45 69.41 69.53 27.49 29.63 23.21 83.61 83.55 83.73 77.78 81.48 70.38 47.33 47.87 46.25 50.86 53.00 46.58 52.29 55.00 46.87
LongCat-Flash-Thinking 3.04 3.10 3.08 3.20 3.18 3.24 30.70 33.90 23.80 68.83 69.54 67.41 69.07 69.31 68.59 31.47 33.83 26.75 80.10 80.62 79.06 78.50 82.46 70.58 47.37 48.69 44.73 53.18 54.45 50.64 53.27 55.06 49.69
GLM-4.6 (w/o thinking) 2.86 3.01 2.58 3.86 4.01 3.57 29.00 29.90 27.20 76.45 76.79 75.77 70.37 70.15 70.81 35.40 38.47 29.26 81.40 81.02 82.16 74.28 78.90 65.04 48.44 49.26 46.80 50.79 53.11 46.15 52.76 55.21 47.86
GLM-4.6 (w/ thinking) 3.08 3.13 2.97 4.06 3.97 3.86 32.80 33.90 30.60 76.83 77.59 75.31 70.27 70.33 70.15 37.48 39.27 33.90 81.30 81.20 81.50 77.66 80.41 72.16 48.90 49.33 48.04 52.67 54.18 49.65 54.43 54.98 53.33
Deepseek-V3.2 (w/o thinking) 3.19 3.16 3.26 4.18 4.15 4.26 31.70 34.50 26.10 77.08 77.19 76.86 70.12 70.26 69.84 35.96 37.68 32.52 81.52 81.61 81.34 75.51 79.80 66.93 48.49 49.41 46.65 52.23 54.76 47.17 54.33 57.02 48.95
Deepseek-V3.2 (w/ thinking) 3.21 3.16 3.31 4.20 4.15 4.30 35.60 38.70 28.90 77.56 77.86 76.96 70.92 70.59 71.58 39.78 38.27 42.80 81.13 81.65 80.09 75.58 79.94 66.86 48.86 49.74 47.10 52.62 54.28 49.30 54.83 57.41 49.67
Appendix B Dataset Construction and Task Design
B.1. LocalSearchBench Task Details
This subsection describes the structure and design principles of the tasks included in LocalSearchBench. Each task is a multi-hop question derived from a merchant record and (optionally) external web evidence; tasks typically require 3–5 reasoning hops to reach a complete, grounded answer. A single task entry consists of four components: (1) the original user query, (2) an explicit hop plan describing the required sub-queries, (3) retrieved evidence snippets used as grounding, and (4) a gold answer annotated with explicit citations to the source records or web pages.
Tasks are categorized by difficulty (L1–L5) according to ambiguity, hop count, and required reasoning steps. The dataset was constructed through an automated generation pipeline followed by a two-stage validation process combining model-based checks and human-in-the-loop verification; the annotation protocol, validation thresholds, and sampling procedures are described in Section 3.3. These design choices aim to ensure that LocalSearchBench covers a broad spectrum of real-world local life service scenarios while maintaining high-quality, verifiable multi-hop reasoning examples.
Table 17. Human–AI agreement rate by field.
Field Human-AI Agreement Rate
subcategory 100.00%
district 100.00%
category 100.00%
city 100.00%
review_count 96.97%
price_range 95.96%
business_hours 95.96%
landmark 92.93%
rating 91.92%
delivery_available 91.82%
delivery_range 91.82%
delivery_fee 87.78%
min_order_amount 86.77%
address 78.79%
phone 74.75%
name 73.74%
mobile 72.72%
products 72.72%
email 69.70%
description 63.64%
Average 88.34%
B.2. Data Validation Evaluation Frameworks
The Data Validation Agent performs a two-stage quality assessment using an LLM-as-a-Judge framework. The evaluation frameworks and validation thresholds are detailed in Table 3.
B.3. Multi-hop QA Examples
Figures 1013 illustrate example multi-hop QA tasks (Hops 3–5, Difficulty: L4). Each figure shows the original user query, the hop-by-hop retrieved evidence and intermediate reasoning steps, and the final assembled answer with citations. Together, these examples demonstrate the system’s chain retrieval and evidence aggregation across regions and domains, and highlight requirements for reasoning about retrieval order, evidence reliability, and quantitative aggregation (e.g., budget calculations) when producing final responses.
Table 18. Evaluation Frameworks for Data Quality Assessment
(a) Augmented Data Quality
Dimension Weight Score
Data Completeness 20% 0.8605
Logical Consistency 30% 0.9000
Geographic Accuracy 10% 0.8568
Content Quality 20% 0.8521
Format Standardization 10% 0.9060
Innovation Degree 10% 0.7080
Overall 100% 0.8596
(b) Privacy Rewriting Data Quality
Dimension Weight Score
Anonymization Degree 50% 0.9473
Logical Consistency 20% 0.8383
Content Consistency 20% 0.9919
Innovation Degree 10% 0.8200
Overall 100% 0.9217
Figure 10. An Example of Multi-Hop QA — Hops: 3 — Difficulty: L4
Figure 11. An Example of Multi-Hop QA — Hops: 4 — Difficulty: L4
Figure 12. An Example of Multi-Hop QA — Hops: 5 — Difficulty: L4
Figure 13. An Example of Multi-Hop QA — Hops: 4 — Difficulty: L4
Figure 14. Prompt for Data Augmentation
Figure 15. Prompt for Privacy Rewriting
Figure 16. Prompt for Augmentation Validation
Input: User query $Q$, Maximum rounds NN
Output: Retrieved information set $\mathcal{I}$
1
// Phase 1: Initialization
$\mathit{information}\leftarrow\emptyset$ ;
// Initialize empty information set
$\mathit{round}\leftarrow 1$ ;
// Initialize iteration counter
2
// Phase 2: Iterative Retrieval Loop
3 for $i\leftarrow 1$ to NN do
4 $\mathit{web\_results}\leftarrow\emptyset$, $\mathit{rag\_results}\leftarrow\emptyset$;
5
 // LLM Response Parsing
6 $\mathit{response}\leftarrow$ LLM($Q$, $\mathit{information}$);
7 $\mathit{web\_query}\leftarrow$ ExtractTag($\mathit{response}$, “<web_search>”);
8 $\mathit{rag\_query}\leftarrow$ ExtractTag($\mathit{response}$, “<rag>”);
9
 // Web Search (at most once per round)
10 if $\mathit{web\_query}\neq\emptyset$ then
11    $\mathit{web\_results}\leftarrow$ WebSearch($\mathit{web\_query}$);
    // Query temporal facts, weather, news, events
12
13
 // LocalRAG Retrieval (at most once per round)
14 if $\mathit{rag\_query}\neq\emptyset$ then
15    $\mathit{rag\_results}\leftarrow$ LocalRAGSearch($\mathit{rag\_query}$);
    // Semantic similarity + geographical proximity analysis
    // Retrieve merchants, POIs, service descriptions
16
17
18 $\mathit{information}\leftarrow\mathit{information}\cup\mathit{web\_results}\cup\mathit{rag\_results}$;
19
20 if SufficientInfo($\mathit{information}$, $Q$) or $i=N$ then
21    break;
22
23
24
// Phase 3: Return Results
25 return $\mathit{information}$;
26
Algorithm 1 Algorithm of Iterative Search